Jefferson Lab HALL A Logbook for February 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Tuesday

  • 35076 02/29/00 23:51 x. Jiang shift activity.
  • 35075 02/29/00 23:51 x. jiang h-arm vernier reading.
  • 35074 02/29/00 23:04 adev Accepted_Events_Run_2476
  • 35071 02/29/00 22:39 adev Accepted_Events_Run_2475
  • 35069 02/29/00 22:12 Saha 3He(e,e'p) in Kin#24
  • 35067 02/29/00 22:01 real beam stripcharts
  • 35066 02/29/00 21:08 adev Accepted_Events_Run_2474
  • 35063 02/29/00 20:52 adev Accepted_Events_Run_2473
  • 35060 02/29/00 20:44 adev Accepted_Events_Run_2472
  • 35058 02/29/00 20:41 x. jiang fpp threshold and chamber voltage. at run 2472.
  • 35056 02/29/00 20:30 adev Accepted_Events_Run_2471
  • 35054 02/29/00 20:26 x. jiang magnet currents.
  • 35053 02/29/00 20:25 x. jiang beam current chat.
  • 35051 02/29/00 20:01 adev Accepted_Events_Run_2470
  • 35049 02/29/00 19:25 adev Accepted_Events_Run_2469
  • 35046 02/29/00 18:10 adev Accepted_Events_Run_2468
  • 35041 02/29/00 16:50 adev Accepted_Events_Run_2467
  • 35039 02/29/00 16:32 adev Accepted_Events_Run_2466
  • 35037 02/29/00 16:17 Saha Shift Summary, 08:00-16:00, Feb 29, 2000
  • 35034 02/29/00 15:55 kuss bizarre target problem
  • 35033 02/29/00 15:49 garibaldi magnet strip chart
  • 35032 02/29/00 15:47 garibaldi beam strip chart
  • 35031 02/29/00 15:47 Saha main control
  • 35030 02/29/00 15:45 Saha OTR screen
  • 35027 02/29/00 15:25 adev Accepted_Events_Run_2464
  • 35024 02/29/00 14:52 adev Accepted_Events_Run_2463
  • 35021 02/29/00 13:18 adev Accepted_Events_Run_2462
  • 35020 02/29/00 11:50 Saha OTR screen after Mott Measurement
  • 35018 02/29/00 11:01 adev Accepted_Events_Run_2461
  • 35016 02/29/00 10:46 adev Accepted_Events_Run_2460
  • 35013 02/29/00 10:35 Saha Hadron Angle Correction. magnet screen
  • 35012 02/29/00 09:28 saha OTR Screen
  • 35011 02/29/00 09:23 adev Accepted_Events_Run_2459
  • 35008 02/29/00 09:10 garibaldi magnets strip chart
  • 35007 02/29/00 09:06 garibaldi magnets strip charts
  • 35006 02/29/00 09:04 garibaldi main control
  • 35005 02/29/00 08:35 garibaldi beam strip chart
  • 35004 02/29/00 08:20 F.Benmokhtar 2 bbu rates for Kin#14 with
  • 35003 02/29/00 08:06 F.Benmokhtar 00:00-08:00 Shift Summary
  • 35001 02/29/00 07:30 adev Accepted_Events_Run_2458
  • 34997 02/29/00 05:56 adev Accepted_Events_Run_2457
  • 34995 02/29/00 04:58 adev Accepted_Events_Run_2456
  • 34992 02/29/00 03:29 adev Accepted_Events_Run_2455
  • 34987 02/29/00 01:55 adev Accepted_Events_Run_2454
  • 34986 02/29/00 01:49 F. Benmokhta PLOTS (rates kin 24)
  • 34985 02/29/00 01:28 F. Benmokhta Number of rates in kinematic 24
  • 34984 02/29/00 01:21 adev Accepted_Events_Run_2453
  • 34981 02/29/00 00:01 Higinbotham 16:00-24:00 Shift Summary
  • Monday

  • 34978 02/28/00 23:48 C.Furget hallA Display
  • 34977 02/28/00 23:47 C.Furget Beam current stripchart
  • 34976 02/28/00 23:43 C.Furget Magnets stripchart
  • 34975 02/28/00 23:43 adev Accepted_Events_Run_2452
  • 34973 02/28/00 23:33 Higinbotham Kin#24 Magnet settings
  • 34971 02/28/00 22:04 kuss ITC502 controller reset
  • 34970 02/28/00 21:33 adev Accepted_Events_Run_2451
  • 34968 02/28/00 21:22 Higinbotham Kinematics # 20:
  • 34967 02/28/00 21:01 Higinbotham Magnet Setting Cross Norm
  • 34966 02/28/00 21:00 Higinbotham Cross Norm for Kin#24
  • 34964 02/28/00 20:42 adev Accepted_Events_Run_2450
  • 34960 02/28/00 20:21 adev Accepted_Events_Run_2449
  • 34957 02/28/00 20:04 adev Accepted_Events_Run_2448
  • 34955 02/28/00 19:50 adev Accepted_Events_Run_2447
  • 34951 02/28/00 19:43 adev Accepted_Events_Run_2446
  • 34949 02/28/00 19:40 adev Accepted_Events_Run_2445
  • 34945 02/28/00 19:24 adev Accepted_Events_Run_2444
  • 34943 02/28/00 19:17 Higinbotham Count Rate Estimate Kin 20
  • 34942 02/28/00 19:16 adev Accepted_Events_Run_2443
  • 34940 02/28/00 18:00 Epstein Count Estimates Kin 13 & 14
  • 34938 02/28/00 17:41 adev Accepted_Events_Run_2442
  • 34936 02/28/00 17:21 adev Accepted_Events_Run_2441
  • 34933 02/28/00 17:04 Higinbotham OTR Check
  • 34932 02/28/00 16:38 C.Furget hall A display
  • 34931 02/28/00 16:23 C. Furget Magnets stripchart
  • 34930 02/28/00 16:03 Saha Shift Summary, 08:00-16:00, Feb 28, 2000
  • 34926 02/28/00 15:41 adev Accepted_Events_Run_2440
  • 34925 02/28/00 15:14 Garibaldi wrong prescale
  • 34923 02/28/00 15:09 adev Accepted_Events_Run_2439
  • 34920 02/28/00 15:08 adev Accepted_Events_Run_2437
  • 34918 02/28/00 15:03 adev Accepted_Events_Run_2437
  • 34915 02/28/00 15:01 real count estimate for kin 20 (total evts needed=6.8M)
  • 34914 02/28/00 15:01 adev Accepted_Events_Run_2436
  • 34912 02/28/00 15:00 adev Accepted_Events_Run_2435
  • 34908 02/28/00 14:50 adev Accepted_Events_Run_2434
  • 34905 02/28/00 14:47 adev Accepted_Events_Run_2433
  • 34902 02/28/00 13:37 adev Accepted_Events_Run_2432
  • 34900 02/28/00 12:35 adev Accepted_Events_Run_2431
  • 34899 02/28/00 12:30 garibaldi kin # 20 counts estimates
  • 34896 02/28/00 11:54 OTR signal OTR screen
  • 34895 02/28/00 11:53 Saha OTR signal
  • 34893 02/28/00 11:10 adev Accepted_Events_Run_2430
  • 34890 02/28/00 10:49 adev Accepted_Events_Run_2429
  • 34888 02/28/00 10:37 Garibaldi magnets strip charts
  • 34887 02/28/00 10:32 Garibaldi beam strip chart
  • 34886 02/28/00 10:27 real Beam stripcharts
  • 34884 02/28/00 08:10 adev Accepted_Events_Run_2428
  • 34882 02/28/00 07:59 Quemener Shift Summary
  • 34879 02/28/00 07:17 Quemener kinematics #18
  • 34878 02/28/00 07:16 adev Accepted_Events_Run_2427
  • 34874 02/28/00 06:57 adev Accepted_Events_Run_2426
  • 34871 02/28/00 06:46 adev Accepted_Events_Run_2425
  • 34870 02/28/00 06:41 adev Accepted_Events_Run_2424
  • 34865 02/28/00 06:13 adev Accepted_Events_Run_2423
  • 34863 02/28/00 06:08 adev Accepted_Events_Run_2422
  • 34861 02/28/00 05:18 adev Accepted_Events_Run_2421
  • 34858 02/28/00 04:40 Adaq beam strip charts
  • 34856 02/28/00 04:19 adev Accepted_Events_Run_2420
  • 34854 02/28/00 03:59 adaq magnets screen
  • 34852 02/28/00 02:53 adev Accepted_Events_Run_2419
  • 34850 02/28/00 02:50 Higinbotham Kin#13 Continued
  • 34848 02/28/00 02:38 adaq magnets screen
  • 34846 02/28/00 02:08 adev Accepted_Events_Run_2418
  • 34843 02/28/00 01:59 adev Accepted_Events_Run_2417
  • 34841 02/28/00 01:43 adev Accepted_Events_Run_2416
  • 34838 02/28/00 01:21 adev Accepted_Events_Run_2415
  • 34836 02/28/00 00:32 adev Accepted_Events_Run_2414
  • Sunday

  • 34832 02/27/00 23:58 Higinbotham Shift Summary
  • 34831 02/27/00 23:55 F.Benmokhtar Number of events in kin 13 and 14 until now
  • 34830 02/27/00 23:42 adev Accepted_Events_Run_2413
  • 34827 02/27/00 23:18 Higinbotham OTR
  • 34826 02/27/00 23:09 adev Accepted_Events_Run_2412
  • 34822 02/27/00 21:53 adev Accepted_Events_Run_2411
  • 34821 02/27/00 21:48 Higinbotham Magnet Strip Charts
  • 34820 02/27/00 21:17 Higinbotham Kin#14 Part 2 Summary
  • 34816 02/27/00 20:49 Higinbotham Spectrometer Settings Kin13
  • 34815 02/27/00 20:48 Higinbotham Kin 14 finished, Back to 13
  • 34814 02/27/00 20:47 adev Accepted_Events_Run_2409
  • 34811 02/27/00 20:11 adev Accepted_Events_Run_2408
  • 34809 02/27/00 20:01 adev Accepted_Events_Run_2407
  • 34808 02/27/00 20:00 x jiang plot of beam current.
  • 34807 02/27/00 19:57 x. jiang plot of magnets of both arm.
  • 34804 02/27/00 19:35 adev Accepted_Events_Run_2406
  • 34801 02/27/00 19:12 Hiiginbotham OTR Check
  • 34800 02/27/00 19:09 adev Accepted_Events_Run_2405
  • 34797 02/27/00 18:46 adev Accepted_Events_Run_2404
  • 34794 02/27/00 18:27 adev Accepted_Events_Run_2403
  • 34791 02/27/00 17:56 adev Accepted_Events_Run_2402
  • 34788 02/27/00 17:48 adev Accepted_Events_Run_2401
  • 34787 02/27/00 17:48 adev Accepted_Events_Run_2401
  • 34783 02/27/00 17:06 adev Accepted_Events_Run_2400
  • 34780 02/27/00 16:38 adev Accepted_Events_Run_2399
  • 34777 02/27/00 16:15 adev Accepted_Events_Run_2398
  • 34776 02/27/00 16:13 kuss Shift summary
  • 34772 02/27/00 15:51 adev Accepted_Events_Run_2397
  • 34769 02/27/00 15:30 adev Accepted_Events_Run_2396
  • 34766 02/27/00 15:01 kuss ESR problems solved
  • 34765 02/27/00 15:01 adev Accepted_Events_Run_2395
  • 34762 02/27/00 14:35 adev Accepted_Events_Run_2394
  • 34759 02/27/00 14:22 kuss ESR alarm
  • 34758 02/27/00 14:10 adev Accepted_Events_Run_2393
  • 34755 02/27/00 13:41 adev Accepted_Events_Run_2392
  • 34752 02/27/00 13:14 adev Accepted_Events_Run_2391
  • 34749 02/27/00 13:08 kuss BPM/current stripchart
  • 34748 02/27/00 13:07 kuss beam energy strip charts
  • 34747 02/27/00 12:49 adev Accepted_Events_Run_2390
  • 34744 02/27/00 12:26 adev Accepted_Events_Run_2389
  • 34741 02/27/00 12:06 adev Accepted_Events_Run_2388
  • 34738 02/27/00 11:42 adev Accepted_Events_Run_2387
  • 34735 02/27/00 11:17 adev Accepted_Events_Run_2386
  • 34732 02/27/00 11:04 kuss OTR
  • 34731 02/27/00 10:53 C. Furget Beam current stripchart
  • 34728 02/27/00 10:47 C. Furget Magnets stripchart
  • 34727 02/27/00 10:46 adev Accepted_Events_Run_2385
  • 34726 02/27/00 10:22 adev Accepted_Events_Run_2384
  • 34723 02/27/00 10:17 kuss trigger dead time
  • 34722 02/27/00 09:55 adev Accepted_Events_Run_2383
  • 34719 02/27/00 09:32 adev Accepted_Events_Run_2382
  • 34716 02/27/00 09:20 C. Furget Display stripchart
  • 34715 02/27/00 09:02 adev Accepted_Events_Run_2381
  • 34712 02/27/00 08:47 kuss beam back
  • 34711 02/27/00 08:31 kuss beam trouble
  • 34710 02/27/00 08:18 adev Accepted_Events_Run_2380
  • 34706 02/27/00 08:04 adev Accepted_Events_Run_2379
  • 34705 02/27/00 08:02 M. Rvachev Shift summary
  • 34702 02/27/00 07:30 adev Accepted_Events_Run_2378
  • 34699 02/27/00 07:14 M. Rvachev Runs 2376, 2377 are junk
  • 34697 02/27/00 07:00 M. Rvachev EQ3 quenched
  • 34696 02/27/00 05:59 adev Accepted_Events_Run_2377
  • 34692 02/27/00 05:42 adev Accepted_Events_Run_2376
  • 34691 02/27/00 05:25 adev Accepted_Events_Run_2375
  • 34688 02/27/00 05:00 adev Accepted_Events_Run_2374
  • 34685 02/27/00 04:36 adev Accepted_Events_Run_2373
  • 34682 02/27/00 04:25 M.Iodice Magnet strip chart
  • 34681 02/27/00 04:16 M.Iodice other strip charts
  • 34679 02/27/00 04:06 M.Iodice Beam strip charts
  • 34678 02/27/00 04:03 M.Iodice OTR screen
  • 34677 02/27/00 04:01 adev Accepted_Events_Run_2372
  • 34676 02/27/00 04:00 M. Rvachev ER1 occasionally disconnects
  • 34674 02/27/00 03:53 M. Rvachev High multiplicity in hadron VDC
  • 34673 02/27/00 03:39 adev Accepted_Events_Run_2371
  • 34670 02/27/00 03:15 adev Accepted_Events_Run_2370
  • 34667 02/27/00 03:11 M. Rvachev Increased current to 113uA
  • 34666 02/27/00 03:07 M. Rvachev Lost some epics signals on hac
  • 34664 02/27/00 02:44 adev Accepted_Events_Run_2369
  • 34662 02/27/00 02:40 adev Accepted_Events_Run_2368
  • 34659 02/27/00 02:14 adev Accepted_Events_Run_2367
  • 34655 02/27/00 01:37 adev Accepted_Events_Run_2366
  • 34654 02/27/00 01:36 adev Accepted_Events_Run_2365
  • 34650 02/27/00 01:11 adev Accepted_Events_Run_2364
  • 34647 02/27/00 00:45 adev Accepted_Events_Run_2363
  • 34643 02/27/00 00:19 adev Accepted_Events_Run_2362
  • 34641 02/27/00 00:14 aniol entry due 14:30 yesterday
  • Saturday

  • 34638 02/26/00 23:49 adev Accepted_Events_Run_2361
  • 34635 02/26/00 23:35 adev Accepted_Events_Run_2360
  • 34633 02/26/00 23:23 Adaq magnets strip charts
  • 34632 02/26/00 23:01 adev Accepted_Events_Run_2359
  • 34627 02/26/00 22:53 adev Accepted_Events_Run_2358
  • 34624 02/26/00 22:43 adev Accepted_Events_Run_2357
  • 34622 02/26/00 22:29 adev Accepted_Events_Run_2356
  • 34618 02/26/00 22:00 adev Accepted_Events_Run_2355
  • 34616 02/26/00 21:57 Higinbotham Spectrometer Setting Kin 14
  • 34615 02/26/00 21:34 aniol/quemen shift summary for kinematic 13
  • 34614 02/26/00 21:16 adev Accepted_Events_Run_2354
  • 34610 02/26/00 20:43 adev Accepted_Events_Run_2353
  • 34608 02/26/00 20:32 adev Accepted_Events_Run_2352
  • 34605 02/26/00 20:23 adev Accepted_Events_Run_2351
  • 34603 02/26/00 20:15 Higinbotham Kin 13 and 14 Simulations
  • 34600 02/26/00 19:56 adev Accepted_Events_Run_2350
  • 34598 02/26/00 19:52 adev Accepted_Events_Run_2349
  • 34595 02/26/00 18:18 adev Accepted_Events_Run_2348
  • 34592 02/26/00 17:40 G. Quemener Display Summary
  • 34591 02/26/00 17:33 Margaziotis Beam problem
  • 34589 02/26/00 17:30 adev Accepted_Events_Run_2347
  • 34586 02/26/00 17:24 adev Accepted_Events_Run_2346
  • 34585 02/26/00 17:22 adev Accepted_Events_Run_2345
  • 34581 02/26/00 17:13 G. Quemener Beam strip charts
  • 34580 02/26/00 17:12 G. Quemener Beam strip charts
  • 34577 02/26/00 16:06 adev Accepted_Events_Run_2344
  • 34575 02/26/00 16:04 F.Benmokhtar Nb of events in kin 1 to 13 until now
  • 34574 02/26/00 16:04 adev Accepted_Events_Run_2343
  • 34573 02/26/00 15:16 adev Accepted_Events_Run_2342
  • 34569 02/26/00 15:03 aniol beam strip charts
  • 34568 02/26/00 14:57 aniol otr screen
  • 34567 02/26/00 14:24 aniol ioc rests and bcm coefficients
  • 34564 02/26/00 14:17 adev Accepted_Events_Run_2341
  • 34562 02/26/00 13:56 aniol strip charts
  • 34561 02/26/00 13:51 Adaq magnet screen
  • 34559 02/26/00 13:36 adev Accepted_Events_Run_2340
  • 34558 02/26/00 13:30 F.Benmokhtar Cuts
  • 34556 02/26/00 13:17 F.Benmokhtar (Sorry) Used cuts in kinematic 13
  • 34555 02/26/00 13:07 `stevensm no control of electron dipole
  • 34554 02/26/00 13:03 F.Benmokhtar 2bbu. rates for Kin 13
  • 34552 02/26/00 11:22 adev Accepted_Events_Run_2339
  • 34550 02/26/00 10:29 adev Accepted_Events_Run_2338
  • 34547 02/26/00 09:22 adev Accepted_Events_Run_2337
  • 34544 02/26/00 08:46 adev Accepted_Events_Run_2336
  • 34541 02/26/00 08:05 M. Rvachev Shift summary
  • 34540 02/26/00 08:04 adev Accepted_Events_Run_2335
  • 34537 02/26/00 07:58 M. Rvachev E_miss plot
  • 34535 02/26/00 06:50 adev Accepted_Events_Run_2334
  • 34533 02/26/00 06:45 C. Furget Beam stripchart
  • 34532 02/26/00 06:10 C. Furget Beam optics
  • 34531 02/26/00 05:29 M. Rvachev Magnets screen
  • 34530 02/26/00 05:27 M. Rvachev OTR screen
  • 34527 02/26/00 05:13 adev Accepted_Events_Run_2333
  • 34525 02/26/00 03:56 M. Rvachev Harp scan with raster
  • 34524 02/26/00 03:50 M. Rvachev OTR screen
  • 34522 02/26/00 03:45 adev Accepted_Events_Run_2332
  • 34520 02/26/00 03:32 C.Furget BeO target spot check
  • 34519 02/26/00 03:14 M. Rvachev Harp scan
  • 34517 02/26/00 03:03 adev Accepted_Events_Run_2331
  • 34514 02/26/00 02:55 adev Accepted_Events_Run_2330
  • 34512 02/26/00 00:20 Higinbotham Shift Summary
  • Friday

  • 34511 02/25/00 20:27 Higinbotham Spectrometer Settings Kin13
  • 34509 02/25/00 19:39 adev Accepted_Events_Run_2328
  • 34507 02/25/00 19:39 adev Accepted_Events_Run_2328
  • 34503 02/25/00 18:38 adev Accepted_Events_Run_2327
  • 34502 02/25/00 18:17 Higinbotham VDC Check
  • 34501 02/25/00 17:28 adev Accepted_Events_Run_2326
  • Thursday

    
    

    Wednesday

  • 34493 02/23/00 21:04 adev Accepted_Events_Run_2325
  • Tuesday

  • 34491 02/22/00 13:24 mitchell ep measurement - update
  • Monday

  • 34484 02/21/00 22:34 adev Accepted_Events_Run_2324
  • 34483 02/21/00 13:20 adev Accepted_Events_Run_2323
  • 34480 02/21/00 09:18 M Jones cosmic run to check fpp
  • Sunday

  • 34477 02/20/00 20:24 folts magnet power down
  • 34476 02/20/00 12:20 mitchell Leaving
  • 34475 02/20/00 12:18 mitchell ep energy measurement
  • 34474 02/20/00 07:19 Eugene Detector tests
  • Saturday

  • 34473 02/19/00 14:04 Suleiman New Plan
  • 34472 02/19/00 13:59 adev Accepted_Events_Run_2322
  • 34471 02/19/00 13:32 suleiman Shifts Cancelled
  • 34470 02/19/00 13:12 kuss target lifter status
  • 34469 02/19/00 09:57 Glashausser target warming up
  • 34468 02/19/00 09:52 Kox "shift" summary 0-8 AM
  • 34466 02/19/00 08:12 Glashausser magnet screen
  • 34464 02/19/00 07:10 kox Target standby parameters
  • 34462 02/19/00 00:22 kuss shift summary
  • Friday

  • 34461 02/18/00 20:58 kuss target lifter problem
  • 34460 02/18/00 19:23 adev Accepted_Events_Run_2321
  • 34458 02/18/00 19:18 Z. Chai E_MISS plot
  • 34456 02/18/00 19:04 M. Rvachev End of Run comments changed
  • 34455 02/18/00 18:51 kuss harp scan with raster, correction
  • 34454 02/18/00 18:47 kuss harp scan with raster
  • 34452 02/18/00 18:39 adev Accepted_Events_Run_2320
  • 34451 02/18/00 18:37 kuss spot
  • 34450 02/18/00 18:32 kuss OTR
  • 34448 02/18/00 18:21 kuss harp scan no raster
  • 34446 02/18/00 18:13 adev Accepted_Events_Run_2319
  • 34444 02/18/00 16:38 adev Accepted_Events_Run_2318
  • 34442 02/18/00 16:21 aniol T1/Q1x ratios
  • 34441 02/18/00 16:10 Glashausser Kin #13 summary-Day 2/18
  • 34438 02/18/00 16:04 Fatiha Number of events in kin 1, 2, 4, 5, 8, 7, (1), 10, 11, 14 and 13...
  • 34437 02/18/00 15:43 adev Accepted_Events_Run_2317
  • 34434 02/18/00 14:23 adev Accepted_Events_Run_2314
  • 34432 02/18/00 14:20 adev Accepted_Events_Run_2316
  • 34428 02/18/00 13:33 glashausser eQ2 polarity
  • 34427 02/18/00 13:11 S.Malov magnet settings
  • 34426 02/18/00 13:09 S.Malov wrong electron Q2 polarity
  • 34423 02/18/00 12:43 adev Accepted_Events_Run_2313
  • 34422 02/18/00 12:32 adev Accepted_Events_Run_2312
  • 34419 02/18/00 12:13 S.Malov beam strip charts
  • 34418 02/18/00 12:12 S.Malov magnet strip charts
  • 34416 02/18/00 11:52 adev Accepted_Events_Run_2311
  • 34414 02/18/00 11:51 S.Malov Magnets settings
  • 34413 02/18/00 11:43 S.Malov OTR screen
  • 34411 02/18/00 11:28 glashausser Access report
  • 34410 02/18/00 11:17 glashausser access for Q2 electron fault
  • 34409 02/18/00 09:41 adev Accepted_Events_Run_2310
  • 34407 02/18/00 09:27 glashausser kin #14 rates
  • 34405 02/18/00 07:51 adev Accepted_Events_Run_2309
  • 34402 02/18/00 06:27 aniol magnet and beam current
  • 34401 02/18/00 06:25 aniol magnet screens
  • 34398 02/18/00 05:49 adev Accepted_Events_Run_2308
  • 34396 02/18/00 05:44 adev Accepted_Events_Run_2306
  • 34395 02/18/00 05:44 adev Accepted_Events_Run_2306
  • 34393 02/18/00 05:44 adev Accepted_Events_Run_2307
  • 34391 02/18/00 05:43 aniol harp scan
  • 34390 02/18/00 05:16 aniol otr screen
  • 34387 02/18/00 04:07 aniol 2303 e_miss
  • 34386 02/18/00 03:30 aniol gas alarm
  • 34384 02/18/00 03:23 adev Accepted_Events_Run_2305
  • 34382 02/18/00 01:59 adev Accepted_Events_Run_2304
  • 34381 02/18/00 01:59 adev Accepted_Events_Run_2304
  • 34377 02/18/00 00:54 aniol RF trips and lower beam current
  • Thursday

  • 34373 02/17/00 23:55 Kox Shift Summary 4-12PM
  • 34372 02/17/00 23:36 adev Accepted_Events_Run_2303
  • 34369 02/17/00 23:23 epstein Kin. 13
  • 34368 02/17/00 23:04 kox OTR for run 2303
  • 34367 02/17/00 22:58 epstein kin. 14 rates
  • 34366 02/17/00 22:02 kox OTR measurement
  • 34365 02/17/00 21:42 adev Accepted_Events_Run_2302
  • 34362 02/17/00 21:14 Ole Hansen Beam optics strip chgarts
  • 34361 02/17/00 21:12 Ole Hansen Magnet stripcharts
  • 34359 02/17/00 19:56 adev Accepted_Events_Run_2301
  • 34356 02/17/00 19:51 adev Accepted_Events_Run_2300
  • 34354 02/17/00 19:40 kox trips at 140 microA
  • 34352 02/17/00 19:17 kox Harp scan results run 547
  • 34351 02/17/00 19:16 adev Accepted_Events_Run_2299
  • 34349 02/17/00 19:06 Kox Harp scan
  • 34347 02/17/00 18:53 adev Accepted_Events_Run_2298
  • 34345 02/17/00 18:15 kox Beam parameters + OTR
  • 34344 02/17/00 17:57 kox Target parameters
  • 34343 02/17/00 17:52 kox Beam Parameters chart
  • 34342 02/17/00 17:50 Kox Magnet setting for kin #13
  • 34341 02/17/00 16:47 kuss Emiss of kin #14
  • 34339 02/17/00 16:39 kuss tc-cor for kin #14
  • 34338 02/17/00 16:34 adev Accepted_Events_Run_2297
  • 34335 02/17/00 15:53 kuss tc_cor of kin #14
  • 34334 02/17/00 15:45 adev Accepted_Events_Run_2296
  • 34330 02/17/00 15:03 adev Accepted_Events_Run_2295
  • 34327 02/17/00 14:37 glashausser Kin #14 summary
  • 34326 02/17/00 14:31 Glashausser Change kinematics
  • 34325 02/17/00 14:19 adev Accepted_Events_Run_2294
  • 34323 02/17/00 13:51 adev Accepted_Events_Run_2293
  • 34320 02/17/00 13:33 S.Malov coda problems run #2278
  • 34319 02/17/00 13:28 adev Accepted_Events_Run_2292
  • 34316 02/17/00 13:05 adev Accepted_Events_Run_2291
  • 34313 02/17/00 12:45 adev Accepted_Events_Run_2290
  • 34310 02/17/00 12:29 adev Accepted_Events_Run_2289
  • 34309 02/17/00 12:27 kuss Emiss for kin #14
  • 34306 02/17/00 12:23 kuss tc_cor at different beam currents for kin #14
  • 34305 02/17/00 12:10 kuss Emiss for kin #11
  • 34304 02/17/00 12:07 adev Accepted_Events_Run_2288
  • 34301 02/17/00 12:01 adev Accepted_Events_Run_2286
  • 34299 02/17/00 11:44 adev Accepted_Events_Run_2287
  • 34295 02/17/00 10:56 kuss Emiss for 2259 (kin #11)
  • 34293 02/17/00 10:48 adev Accepted_Events_Run_2285
  • 34291 02/17/00 10:47 S.Malov OTR screen
  • 34290 02/17/00 10:30 adev Accepted_Events_Run_2284
  • 34287 02/17/00 10:13 kuss coincidences in run 2275 (kin #14)
  • 34286 02/17/00 10:09 adev Accepted_Events_Run_2283
  • 34283 02/17/00 09:59 S.Malov beam stripchart
  • 34282 02/17/00 09:58 S.Malov magnet stripchart
  • 34281 02/17/00 09:56 S.Malov magnet settings
  • 34278 02/17/00 09:45 adev Accepted_Events_Run_2282
  • 34277 02/17/00 09:35 adev Accepted_Events_Run_2278
  • 34276 02/17/00 09:35 adev Accepted_Events_Run_2278
  • 34275 02/17/00 09:34 adev Accepted_Events_Run_2278
  • 34274 02/17/00 09:23 adev Accepted_Events_Run_2281
  • 34269 02/17/00 09:17 adev Accepted_Events_Run_2280
  • 34262 02/17/00 08:55 adev Accepted_Events_Run_2277
  • 34257 02/17/00 08:31 adev Accepted_Events_Run_2276
  • 34256 02/17/00 08:13 adev Accepted_Events_Run_2275
  • 34254 02/17/00 08:07 aniol target warmed to 7K
  • 34251 02/17/00 07:51 Fatiha Number of events in kin 1, 2, 4, 5, 8, 7, (1), 10 and 11.
  • 34250 02/17/00 07:48 adev Accepted_Events_Run_2274
  • 34249 02/17/00 07:46 aniol change to kin#14
  • 34245 02/17/00 07:25 adev Accepted_Events_Run_2273
  • 34242 02/17/00 07:20 adev Accepted_Events_Run_2272
  • 34239 02/17/00 07:08 Adaq raster size
  • 34238 02/17/00 06:31 adaq harp #546
  • 34236 02/17/00 06:26 adev Accepted_Events_Run_2271
  • 34234 02/17/00 06:17 Adaq harp 545 results
  • 34233 02/17/00 06:02 Adaq harp scan #545
  • 34231 02/17/00 06:00 adev Accepted_Events_Run_2270
  • 34229 02/17/00 05:45 adev Accepted_Events_Run_2269
  • 34228 02/17/00 05:44 aniol List of runs for kin 11
  • 34225 02/17/00 05:23 adev Accepted_Events_Run_2268
  • 34223 02/17/00 04:58 Z. Chai H-Arm S2-3-L ADC
  • 34222 02/17/00 04:50 adev Accepted_Events_Run_2267
  • 34219 02/17/00 04:34 Adaq otr screen
  • 34218 02/17/00 04:26 aniol strip chart
  • 34217 02/17/00 04:22 adev Accepted_Events_Run_2266
  • 34216 02/17/00 04:22 Adaq strip charts
  • 34212 02/17/00 03:51 adev Accepted_Events_Run_2265
  • 34210 02/17/00 03:39 adev Accepted_Events_Run_2264
  • 34206 02/17/00 03:07 adev Accepted_Events_Run_2263
  • 34205 02/17/00 02:51 Z. Chai DAQ problem happened again
  • 34204 02/17/00 02:41 adev Accepted_Events_Run_2262
  • 34202 02/17/00 02:37 Adaq daq problem
  • 34200 02/17/00 02:25 adev Accepted_Events_Run_2261
  • 34197 02/17/00 01:57 adev Accepted_Events_Run_2260
  • 34194 02/17/00 01:32 Adaq anomalous dead time r2259
  • 34193 02/17/00 01:27 adev Accepted_Events_Run_2259
  • 34190 02/17/00 00:58 adev Accepted_Events_Run_2258
  • 34186 02/17/00 00:31 adev Accepted_Events_Run_2257
  • 34184 02/17/00 00:12 LeRose Shift Summary
  • 34182 02/17/00 00:02 adev Accepted_Events_Run_2256
  • Wednesday

  • 34179 02/16/00 23:38 adev Accepted_Events_Run_2255
  • 34176 02/16/00 23:06 adev Accepted_Events_Run_2254
  • 34173 02/16/00 22:38 adev Accepted_Events_Run_2253
  • 34170 02/16/00 22:10 LeRose OTR out
  • 34169 02/16/00 22:08 adev Accepted_Events_Run_2252
  • 34165 02/16/00 21:34 adev Accepted_Events_Run_2251
  • 34162 02/16/00 21:27 adev Accepted_Events_Run_2250
  • 34160 02/16/00 21:19 LeRose Beam Current
  • 34159 02/16/00 21:07 adev Accepted_Events_Run_2249
  • 34156 02/16/00 20:41 LeRose Beam Optics Strip Chart
  • 34155 02/16/00 20:39 LeRose Ytgt Problem: Runs 2208-2244
  • 34154 02/16/00 20:33 LeRose Magnet Strip Charts
  • 34152 02/16/00 20:28 adev Accepted_Events_Run_2248
  • 34149 02/16/00 20:25 Adaq y_tg problem
  • 34148 02/16/00 20:18 LeRose Ytgt Problem solved
  • 34146 02/16/00 20:14 adev Accepted_Events_Run_2247
  • 34144 02/16/00 19:59 LeRose Q2 Electron
  • 34143 02/16/00 18:52 LeRose Ytgt: Q2 Electron Problem
  • 34141 02/16/00 18:33 adev Accepted_Events_Run_2246
  • 34139 02/16/00 18:23 adev Accepted_Events_Run_2245
  • 34136 02/16/00 18:15 adev Accepted_Events_Run_2244
  • 34135 02/16/00 18:13 Nilanga y_tg problem
  • 34131 02/16/00 17:36 adev Accepted_Events_Run_2243
  • 34130 02/16/00 16:55 LeRose ROC 14
  • 34128 02/16/00 16:39 adev Accepted_Events_Run_2241
  • 34127 02/16/00 16:34 Epstein y_tgte spectra
  • 34124 02/16/00 16:30 Epstein Shift Summary
  • 34121 02/16/00 15:34 adev Accepted_Events_Run_2240
  • 34119 02/16/00 15:33 kox Harp scan II
  • 34118 02/16/00 15:27 kox Harp scan
  • 34116 02/16/00 13:35 M Jones changed epicslogger
  • 34115 02/16/00 13:15 adev Accepted_Events_Run_2239
  • 34113 02/16/00 12:59 adev Accepted_Events_Run_2238
  • 34110 02/16/00 12:21 adev Accepted_Events_Run_2237
  • 34107 02/16/00 12:14 Saha MCC Meeting - Higher beam currents
  • 34106 02/16/00 11:40 adev Accepted_Events_Run_2236
  • 34103 02/16/00 10:52 adev Accepted_Events_Run_2235
  • 34099 02/16/00 10:13 M JOnes change epicslogger script
  • 34098 02/16/00 10:02 adev Accepted_Events_Run_2234
  • 34097 02/16/00 10:01 S.Malov magnet strip chart
  • 34095 02/16/00 09:49 Epstein OTR drifts
  • 34094 02/16/00 09:48 kox magnet chart kin#11
  • 34093 02/16/00 09:47 kox beam conditions
  • 34091 02/16/00 09:44 Epstein OTRT
  • 34090 02/16/00 09:28 kox Target chart
  • 34089 02/16/00 09:23 adev Accepted_Events_Run_2233
  • 34087 02/16/00 09:17 Adaq OTR
  • 34085 02/16/00 09:12 adev Accepted_Events_Run_2232
  • 34082 02/16/00 09:01 adev Accepted_Events_Run_2231
  • 34080 02/16/00 08:26 adev Accepted_Events_Run_2230
  • 34076 02/16/00 08:10 M. Rvachev Shift summary
  • 34075 02/16/00 08:07 boeglinw strip chart beam energy and p_e
  • 34074 02/16/00 07:53 adev Accepted_Events_Run_2229
  • 34072 02/16/00 07:52 M. Rvachev OTR screen
  • 34069 02/16/00 07:10 M. Rvachev Changed to kin 11
  • 34068 02/16/00 07:08 adev Accepted_Events_Run_2228
  • 34065 02/16/00 06:46 adev Accepted_Events_Run_2227
  • 34063 02/16/00 06:42 M. Rvachev Run summary for Kin#10
  • 34062 02/16/00 06:38 adev Accepted_Events_Run_2226
  • 34059 02/16/00 06:16 M. Rvachev Magnets screen
  • 34056 02/16/00 04:38 adev Accepted_Events_Run_2225
  • 34055 02/16/00 03:02 adev Accepted_Events_Run_2224
  • 34052 02/16/00 01:44 adev Accepted_Events_Run_2223
  • 34049 02/16/00 01:07 M. Rvachev OTR screen
  • 34048 02/16/00 00:29 adev Accepted_Events_Run_2222
  • Tuesday

  • 34044 02/15/00 23:30 perdrisat optics strip chart
  • 34043 02/15/00 23:19 adev Accepted_Events_Run_2221
  • 34040 02/15/00 22:07 adev Accepted_Events_Run_2220
  • 34039 02/15/00 22:06 perdrisat strip charts
  • 34038 02/15/00 22:03 perdrisat OTR
  • 34035 02/15/00 21:28 Ole Hansen Magnet screen
  • 34034 02/15/00 20:45 perdrisat data, finally
  • 34032 02/15/00 20:23 perdrisat harp pictrure
  • 34031 02/15/00 20:13 perdrisat new harp and Moeller quad situation
  • 34030 02/15/00 19:27 perdrisat spot with harp
  • 34029 02/15/00 19:26 kuss spot for 2219
  • 34028 02/15/00 19:25 adev Accepted_Events_Run_2219
  • 34025 02/15/00 19:16 perdrisat final harp
  • 34024 02/15/00 17:58 perdrisat harp
  • 34023 02/15/00 17:17 perdrisat OTR again
  • 34021 02/15/00 16:48 adev Accepted_Events_Run_2218
  • 34019 02/15/00 16:36 perdrisat OTr scan begin of shift
  • 34018 02/15/00 16:03 S. Strauch Shift Summary (part of kin. #10 run summary)
  • 34015 02/15/00 08:00 adev Accepted_Events_Run_2217
  • 34013 02/15/00 07:54 Eric Voutier Shift Summary
  • 34012 02/15/00 07:53 W. Boeglin counts in Emiss for kin 10
  • 34007 02/15/00 05:57 adev Accepted_Events_Run_2215
  • 34006 02/15/00 05:06 adev Accepted_Events_Run_2214
  • 34002 02/15/00 03:41 adev Accepted_Events_Run_2212
  • 33999 02/15/00 02:02 Eric Voutier HARP scan
  • 33998 02/15/00 01:49 Eric Voutier Beam parameters
  • 33997 02/15/00 01:42 adev Accepted_Events_Run_2211
  • 33995 02/15/00 01:12 Kox Beam chart
  • 33994 02/15/00 01:11 kox Magnet chart
  • 33993 02/15/00 00:56 Kox target chart
  • 33992 02/15/00 00:25 adev Accepted_Events_Run_2210
  • Monday

  • 33989 02/14/00 23:57 Ole Hansen Shift Summary
  • 33987 02/14/00 23:18 M. Rvachev reactez plot
  • 33986 02/14/00 23:06 Ole Hansen Current reduced to 98uA
  • 33984 02/14/00 23:05 Ole Hansen reactez resolution deteriorates
  • 33983 02/14/00 23:04 Ole Hansen OTR screen
  • 33982 02/14/00 23:03 adev Accepted_Events_Run_2209
  • 33980 02/14/00 22:36 liatard main control display
  • 33979 02/14/00 22:33 Liatard beam stripchart
  • 33978 02/14/00 21:45 adev Accepted_Events_Run_2208
  • 33975 02/14/00 20:44 Ole Hansen Target vacuum high but stable
  • 33974 02/14/00 20:40 Ole Hansen OTR screen
  • 33973 02/14/00 20:35 adev Accepted_Events_Run_2207
  • 33970 02/14/00 20:10 Ole Hansen Target vacuum rising
  • 33969 02/14/00 19:24 Ole Hansen Start data taking
  • 33968 02/14/00 19:24 Ole Hansen Mx. beam current is 110uA for now
  • 33966 02/14/00 19:08 adev Accepted_Events_Run_2206
  • 33964 02/14/00 19:07 Ole Hansen OTR screen at 50uA
  • 33963 02/14/00 19:00 mitchell ep energy measurement online result
  • 33960 02/14/00 16:03 J. P. Chen shift summary
  • 33959 02/14/00 15:53 S. Strauch Magnet Stripchart
  • 33958 02/14/00 15:16 adev Accepted_Events_Run_2205
  • 33954 02/14/00 14:54 adev Accepted_Events_Run_2204
  • 33952 02/14/00 14:42 Saha Hall A Beam Optics studies
  • 33951 02/14/00 13:39 Saha Arc Energy Measurment, Feb 14, 2000, 13:00
  • 33950 02/14/00 10:20 adev Accepted_Events_Run_2203
  • 33948 02/14/00 10:12 adev Accepted_Events_Run_2202
  • 33945 02/14/00 09:47 adev Accepted_Events_Run_2201
  • 33942 02/14/00 09:23 adev Accepted_Events_Run_2200
  • 33938 02/14/00 08:49 adev Accepted_Events_Run_2199
  • 33935 02/14/00 08:45 kuss target pressure
  • 33934 02/14/00 08:34 S. Strauch Megnet Screen for Kinematics #1
  • 33933 02/14/00 08:21 kuss shift summary
  • 33931 02/14/00 08:03 kuss OTR
  • 33930 02/14/00 07:59 adev Accepted_Events_Run_2198
  • 33928 02/14/00 07:43 fatiha Number of events in kin 1, 2, 4, 5, 8, 7
  • 33927 02/14/00 06:52 kuss ITC502_1 LED display does not work
  • 33926 02/14/00 06:40 kuss beam current stripchart
  • 33925 02/14/00 06:39 kuss beam optics stripchart
  • 33924 02/14/00 06:36 adev Accepted_Events_Run_2197
  • 33923 02/14/00 06:35 kuss hall A status screen
  • 33920 02/14/00 06:33 kuss magnet stripchart
  • 33919 02/14/00 06:32 kuss beam current stripchart
  • 33918 02/14/00 05:22 kuss OTR
  • 33917 02/14/00 05:08 kuss beam current
  • 33915 02/14/00 05:03 kuss OTR
  • 33914 02/14/00 05:02 adev Accepted_Events_Run_2196
  • 33912 02/14/00 04:21 fatiha 2bbu rates for kin # 7
  • 33911 02/14/00 03:41 adev Accepted_Events_Run_2195
  • 33908 02/14/00 02:50 kuss beam current
  • 33907 02/14/00 02:49 kuss magnet stripcharts
  • 33906 02/14/00 02:47 kuss beam current
  • 33905 02/14/00 02:34 kuss beam optics strip charts
  • 33904 02/14/00 02:27 kuss Magnet stripchart
  • 33903 02/14/00 02:21 kuss OTR
  • 33902 02/14/00 02:20 kuss OTR
  • 33901 02/14/00 02:16 adev Accepted_Events_Run_2194
  • 33898 02/14/00 02:11 kuss new hdr_he3_7
  • 33897 02/14/00 02:07 kuss target IOC reboot
  • 33896 02/14/00 01:24 Fatiha Cerenkov ADCs
  • 33893 02/14/00 00:46 adev Accepted_Events_Run_2193
  • 33891 02/14/00 00:03 Kox Shift Summary 4-12 PM
  • Sunday

  • 33890 02/13/00 23:33 adev Accepted_Events_Run_2192
  • 33887 02/13/00 23:23 Nilanga OTR
  • 33885 02/13/00 22:08 adev Accepted_Events_Run_2191
  • 33882 02/13/00 21:00 kox beam conditions
  • 33881 02/13/00 20:57 kox OTR measurement
  • 33880 02/13/00 20:47 adev Accepted_Events_Run_2190
  • 33878 02/13/00 19:42 kox beam pos. chart
  • 33876 02/13/00 18:34 kox Target chart
  • 33875 02/13/00 18:31 kox Magnet seetings + charts
  • 33873 02/13/00 18:21 Kox OTR + beam conditions
  • 33872 02/13/00 18:06 Kox Harp scan
  • 33871 02/13/00 17:25 adev Accepted_Events_Run_2188
  • 33869 02/13/00 16:27 Epstein OTR
  • 33868 02/13/00 16:18 adev Accepted_Events_Run_2187
  • 33864 02/13/00 15:58 M. Rvachev Shift summary
  • 33863 02/13/00 15:27 M. Rvachev OTR
  • 33862 02/13/00 14:58 Liatard magnet stripcharts
  • 33861 02/13/00 14:56 Liatard beam stripchart
  • 33859 02/13/00 14:34 adev Accepted_Events_Run_2186
  • 33857 02/13/00 14:24 M. Rvachev New espace kumac available
  • 33855 02/13/00 14:06 adev Accepted_Events_Run_2185
  • 33852 02/13/00 13:40 Higinbotham OTR
  • 33851 02/13/00 13:30 adev Accepted_Events_Run_2184
  • 33849 02/13/00 12:10 M. Rvachev OTR for run 2184
  • 33847 02/13/00 12:07 adev Accepted_Events_Run_2183
  • 33845 02/13/00 12:06 kuss analysis harp scan kin #7
  • 33843 02/13/00 11:42 kuss harp scan for kin #7
  • 33842 02/13/00 11:41 adev Accepted_Events_Run_2182
  • 33839 02/13/00 11:33 M. Rvachev Changed to kin 7
  • 33838 02/13/00 11:08 adev Accepted_Events_Run_2181
  • 33836 02/13/00 11:03 Nilanga Counts from #8
  • 33834 02/13/00 10:47 adev Accepted_Events_Run_2180
  • 33831 02/13/00 09:41 adev Accepted_Events_Run_2179
  • 33829 02/13/00 08:59 Liatard magnet stripcharts
  • 33828 02/13/00 08:57 adev Accepted_Events_Run_2178
  • 33827 02/13/00 08:54 Liatard beam stripchart
  • 33824 02/13/00 08:01 R. Michaels owl shift summary
  • 33822 02/13/00 07:53 R. Michaels spot++
  • 33821 02/13/00 07:51 adev Accepted_Events_Run_2177
  • 33818 02/13/00 07:10 W.Boeglin, R correction on statistic estimate
  • 33817 02/13/00 06:59 W. Boeglin/ statistics for kinematics 8
  • 33816 02/13/00 06:56 adev Accepted_Events_Run_2176
  • 33813 02/13/00 05:58 adev Accepted_Events_Run_2175
  • 33810 02/13/00 05:04 adev Accepted_Events_Run_2174
  • 33807 02/13/00 04:10 R. Michaels magnets, epics screen
  • 33806 02/13/00 04:09 R. Michaels spect_p0
  • 33805 02/13/00 04:08 R. Michaels beam strip chart
  • 33804 02/13/00 04:06 R. Michaels OTR screen
  • 33803 02/13/00 04:06 adev Accepted_Events_Run_2173
  • 33800 02/13/00 03:11 adev Accepted_Events_Run_2172
  • 33797 02/13/00 02:23 adev Accepted_Events_Run_2171
  • 33794 02/13/00 01:42 Eric Voutier Target Chart
  • 33793 02/13/00 01:21 adev Accepted_Events_Run_2170
  • 33790 02/13/00 00:27 adev Accepted_Events_Run_2169
  • Saturday

  • 33785 02/12/00 23:28 adev Accepted_Events_Run_2168
  • 33782 02/12/00 23:15 Tieulent HARP scan
  • 33781 02/12/00 23:05 Tieulent Spot ++
  • 33779 02/12/00 23:04 adev Accepted_Events_Run_2167
  • 33777 02/12/00 22:12 Tieulent spectrometer setting kin 8
  • 33776 02/12/00 21:50 adev Accepted_Events_Run_2166
  • 33774 02/12/00 21:32 adev Accepted_Events_Run_2165
  • 33771 02/12/00 21:18 adev Accepted_Events_Run_2164
  • 33767 02/12/00 20:43 Marty Epstei Finished Kinematics 5
  • 33766 02/12/00 20:41 adev Accepted_Events_Run_2163
  • 33764 02/12/00 20:11 Adaq EARM dipole drift
  • 33763 02/12/00 20:03 Tieulent OTR OUT
  • 33762 02/12/00 19:59 adev Accepted_Events_Run_2162
  • 33759 02/12/00 19:39 Marty Epstei Kin. 5 count rate
  • 33757 02/12/00 19:07 adev Accepted_Events_Run_2161
  • 33755 02/12/00 19:02 Tieulent otr
  • 33754 02/12/00 18:24 adev Accepted_Events_Run_2160
  • 33751 02/12/00 17:57 Tieulent Beam parameters
  • 33750 02/12/00 17:55 Tieulent Spectrometers strip charts
  • 33749 02/12/00 17:51 Adaq Kin4: MCEEP vs. data
  • 33747 02/12/00 17:20 kuss restore beam for data taking
  • 33746 02/12/00 17:19 kuss current scan 10,50,100#&181;A
  • 33744 02/12/00 16:59 adev Accepted_Events_Run_2159
  • 33742 02/12/00 16:43 adev Accepted_Events_Run_2158
  • 33739 02/12/00 16:36 adev Accepted_Events_Run_2157
  • 33736 02/12/00 16:31 adev Accepted_Events_Run_2156
  • 33733 02/12/00 16:26 adev Accepted_Events_Run_2155
  • 33730 02/12/00 16:16 adev Accepted_Events_Run_2154
  • 33727 02/12/00 16:09 adev Accepted_Events_Run_2153
  • 33724 02/12/00 16:03 kuss spot of 2152
  • 33723 02/12/00 16:02 adev Accepted_Events_Run_2152
  • 33720 02/12/00 15:58 M. Rvachev Shift summary
  • 33718 02/12/00 15:16 M. Rvachev Changed to kin 5
  • 33717 02/12/00 15:01 adev Accepted_Events_Run_2151
  • 33714 02/12/00 14:40 adev Accepted_Events_Run_2150
  • 33713 02/12/00 14:39 M. Rvachev List of runs for kin 4
  • 33710 02/12/00 13:29 adev Accepted_Events_Run_2149
  • 33707 02/12/00 13:17 adev Accepted_Events_Run_2148
  • 33704 02/12/00 13:04 adev Accepted_Events_Run_2147
  • 33701 02/12/00 12:51 adev Accepted_Events_Run_2146
  • 33699 02/12/00 12:41 M. Rvachev List of dummy runs for kin 1
  • 33698 02/12/00 12:35 M. Rvachev omega-q cuts for kin 4
  • 33697 02/12/00 12:06 adev Accepted_Events_Run_2145
  • 33694 02/12/00 11:07 adev Accepted_Events_Run_2144
  • 33691 02/12/00 11:01 M. Rvachev OTR for run 2144
  • 33690 02/12/00 10:20 kox Target chart begining Kin #4
  • 33688 02/12/00 10:12 adev Accepted_Events_Run_2143
  • 33685 02/12/00 10:03 adev Accepted_Events_Run_2142
  • 33682 02/12/00 09:55 adev Accepted_Events_Run_2141
  • 33679 02/12/00 09:43 adev Accepted_Events_Run_2140
  • 33677 02/12/00 09:39 M. Rvachev Changed to kin 4
  • 33676 02/12/00 09:24 Higinbotham 2 Body Counts In Kin#1,2
  • 33674 02/12/00 09:09 adev Accepted_Events_Run_2139
  • 33671 02/12/00 08:54 adev Accepted_Events_Run_2138
  • 33668 02/12/00 08:44 adev Accepted_Events_Run_2137
  • 33666 02/12/00 08:40 M. Rvachev Beam parameters and mignet screen
  • 33664 02/12/00 08:28 adev Accepted_Events_Run_2136
  • 33662 02/12/00 08:26 adev Accepted_Events_Run_2135
  • 33658 02/12/00 08:02 fatiha Number of events during the shift
  • 33657 02/12/00 08:01 R. Michaels owl shift summary
  • 33654 02/12/00 07:56 adev Accepted_Events_Run_2134
  • 33651 02/12/00 07:53 adev Accepted_Events_Run_2133
  • 33649 02/12/00 07:49 adev Accepted_Events_Run_2132
  • 33646 02/12/00 07:10 adev Accepted_Events_Run_2131
  • 33643 02/12/00 06:44 R. Michaels OTR background
  • 33641 02/12/00 06:38 adev Accepted_Events_Run_2130
  • 33640 02/12/00 06:33 adev Accepted_Events_Run_2129
  • 33636 02/12/00 05:39 adev Accepted_Events_Run_2128
  • 33633 02/12/00 05:01 R. Michaels 6 msr collimator inserted
  • 33632 02/12/00 05:01 adev Accepted_Events_Run_2127
  • 33629 02/12/00 04:57 fatiha Rates for kinematic 1
  • 33628 02/12/00 04:54 adev Accepted_Events_Run_2126
  • 33626 02/12/00 04:25 adev Accepted_Events_Run_2125
  • 33623 02/12/00 04:04 R. Michaels cpu states when ER1/deadtime problems
  • 33622 02/12/00 03:54 adev Accepted_Events_Run_2124
  • 33619 02/12/00 03:32 R. Michaels field regulation ?
  • 33618 02/12/00 03:23 adev Accepted_Events_Run_2123
  • 33615 02/12/00 02:59 R. Michaels spect_p0
  • 33614 02/12/00 02:58 R. Michaels beam strip chart
  • 33613 02/12/00 02:56 adev Accepted_Events_Run_2122
  • 33610 02/12/00 02:39 adev Accepted_Events_Run_2121
  • 33607 02/12/00 02:12 adev Accepted_Events_Run_2120
  • 33604 02/12/00 01:52 fatiha Target chart 2120
  • 33603 02/12/00 01:45 R. Michaels OTR screen, 100 uA beam
  • 33601 02/12/00 01:31 adev Accepted_Events_Run_2119
  • 33599 02/12/00 01:18 R. Michaels checks of BPM and Raster
  • 33598 02/12/00 01:11 fatiha e_miss spectra
  • 33597 02/12/00 01:04 adev Accepted_Events_Run_2118
  • 33594 02/12/00 00:57 R. Michaels EPICS, magnets
  • 33591 02/12/00 00:35 adev Accepted_Events_Run_2116
  • 33587 02/12/00 00:13 adev Accepted_Events_Run_2115
  • 33586 02/12/00 00:12 R. Michaels beam ok, start kinematics 1,2
  • 33583 02/12/00 00:06 adev Accepted_Events_Run_2114
  • Friday

  • 33580 02/11/00 23:59 perdrisat more harp scans
  • 33578 02/11/00 23:50 adev Accepted_Events_Run_2113
  • 33576 02/11/00 23:00 perdrisat harp scan galore
  • 33574 02/11/00 22:36 adev Accepted_Events_Run_2112
  • 33572 02/11/00 22:34 adev Accepted_Events_Run_2111
  • 33569 02/11/00 22:30 adev Accepted_Events_Run_2110
  • 33565 02/11/00 22:16 adev Accepted_Events_Run_2109
  • 33564 02/11/00 21:34 perdrisat beam energy display
  • 33562 02/11/00 21:21 adev Accepted_Events_Run_2108
  • 33560 02/11/00 20:28 perdrisat raster+spot problems
  • 33558 02/11/00 18:17 adev Accepted_Events_Run_2107
  • 33556 02/11/00 18:16 adev Accepted_Events_Run_2106
  • 33552 02/11/00 18:13 adev Accepted_Events_Run_2105
  • 33549 02/11/00 18:04 adev Accepted_Events_Run_2104
  • 33545 02/11/00 17:39 adev Accepted_Events_Run_2103
  • 33543 02/11/00 16:46 adev Accepted_Events_Run_2102
  • 33539 02/11/00 15:59 Tieulent Shift summary
  • 33538 02/11/00 15:39 adev Accepted_Events_Run_2101
  • 33536 02/11/00 15:35 adev Accepted_Events_Run_2100
  • 33531 02/11/00 15:07 adev Accepted_Events_Run_2099
  • 33529 02/11/00 14:59 Tieulent Target Parameters
  • 33528 02/11/00 14:57 Tieulent target parameters
  • 33527 02/11/00 14:50 Saha Beam Position and size
  • 33525 02/11/00 14:43 adev Accepted_Events_Run_2098
  • 33524 02/11/00 14:39 S.Malov beam spot
  • 33523 02/11/00 14:38 Saha Harp Scan - Raster ON
  • 33520 02/11/00 14:33 adev Accepted_Events_Run_2097
  • 33517 02/11/00 14:13 adev Accepted_Events_Run_2096
  • 33515 02/11/00 14:11 S.Malov beam stripchart
  • 33513 02/11/00 14:08 S.Malov magnets status
  • 33511 02/11/00 12:04 adev Accepted_Events_Run_2095
  • 33509 02/11/00 11:51 adev Accepted_Events_Run_2094
  • 33506 02/11/00 11:30 tieulent otr
  • 33505 02/11/00 10:50 adaq
  • 33504 02/11/00 10:41 Marty Cuts for 4.8 GeV running
  • 33503 02/11/00 10:13 tieulent Spectrometer strip chart
  • 33502 02/11/00 10:12 tieulent e89044_beam
  • 33501 02/11/00 08:55 tieulent Spectrometer setting kin#1
  • 33500 02/11/00 08:46 adev Accepted_Events_Run_2093
  • 33498 02/11/00 08:35 Kox Shift Summary 0-8 AM
  • 33494 02/11/00 08:01 Kox harp scan
  • 33493 02/11/00 07:58 R. Michaels Beam size from BPMs (Harp Scan)
  • 33491 02/11/00 07:52 adev Accepted_Events_Run_2092
  • 33488 02/11/00 07:45 adev Accepted_Events_Run_2091
  • 33486 02/11/00 07:38 adev Accepted_Events_Run_2090
  • 33483 02/11/00 06:15 boeglin kinematic correction
  • 33482 02/11/00 06:11 adev Accepted_Events_Run_2089
  • 33480 02/11/00 04:52 R. Michaels download trig, change E-arm momentum
  • 33475 02/11/00 04:29 R. Michaels magnet epics screen
  • 33474 02/11/00 04:28 R. Michaels spect_p0
  • 33473 02/11/00 04:27 R. Michaels beam strip chart
  • 33472 02/11/00 02:39 R. Michaels He3(e,e') singles runs
  • 33471 02/11/00 02:33 adev Accepted_Events_Run_2087
  • 33467 02/11/00 02:17 adev Accepted_Events_Run_2086
  • 33465 02/11/00 02:15 adev Accepted_Events_Run_2085
  • 33463 02/11/00 01:49 adev Accepted_Events_Run_2084
  • 33458 02/11/00 01:17 adev Accepted_Events_Run_2083
  • 33455 02/11/00 01:09 adev Accepted_Events_Run_2082
  • 33453 02/11/00 01:04 mitchell ep energy measurement
  • Thursday

  • 33452 02/10/00 23:58 Ole Hansen Shift Summary
  • 33450 02/10/00 23:43 Saha Arc Energy Measurement, Feb 10,2000
  • 33449 02/10/00 23:40 Ole Hansen Beam strip chart
  • 33448 02/10/00 22:09 Akila current calibration
  • 33447 02/10/00 21:48 kuss beam strip chanrt
  • 33446 02/10/00 21:47 kuss current calibration #2
  • 33445 02/10/00 21:46 adev Accepted_Events_Run_2081
  • 33443 02/10/00 21:30 kuss Emiss resolution for kin #17 (Dec 1999)
  • 33442 02/10/00 21:11 kuss HRSH settings change
  • 33440 02/10/00 21:09 kuss current calibration, step down
  • 33439 02/10/00 21:07 adev Accepted_Events_Run_2080
  • 33436 02/10/00 20:06 adev Accepted_Events_Run_2079
  • 33433 02/10/00 19:47 kuss magnet striptool for kin #36
  • 33432 02/10/00 19:44 adev Accepted_Events_Run_2078
  • 33430 02/10/00 19:38 Ole Hansen Elastic numbers look good to us
  • 33429 02/10/00 19:33 adev Accepted_Events_Run_2077
  • 33426 02/10/00 19:19 adev Accepted_Events_Run_2076
  • 33423 02/10/00 19:07 adev Accepted_Events_Run_2075
  • 33420 02/10/00 18:54 adev Accepted_Events_Run_2074
  • 33417 02/10/00 18:51 Ole Hansen Run 2072 might be useful after all
  • 33416 02/10/00 18:46 Ole Hansen Run 2073 is junk
  • 33415 02/10/00 18:44 Ole Hansen Wrong start-of-run header for run 2073
  • 33414 02/10/00 18:44 adev Accepted_Events_Run_2073
  • 33412 02/10/00 18:39 Ole Hansen Ignore run 2072
  • 33409 02/10/00 18:07 adev Accepted_Events_Run_2072
  • 33407 02/10/00 18:01 Akila magnet screen
  • 33405 02/10/00 17:56 adev Accepted_Events_Run_2071
  • 33402 02/10/00 16:52 adev Accepted_Events_Run_2070
  • 33400 02/10/00 16:38 Akila magnetic screen
  • 33399 02/10/00 16:28 A. Saha Shift Summary
  • 33397 02/10/00 16:25 adev Accepted_Events_Run_2069
  • 33393 02/10/00 16:08 adev Accepted_Events_Run_2068
  • 33390 02/10/00 15:50 adev Accepted_Events_Run_2067
  • 33387 02/10/00 15:47 adev Accepted_Events_Run_2066
  • 33385 02/10/00 15:46 adev Accepted_Events_Run_2065
  • 33382 02/10/00 15:34 S.Malov magnet settings
  • 33381 02/10/00 15:31 S.Malov beam spot
  • 33379 02/10/00 15:24 Eric Voutier Harp Scan
  • 33378 02/10/00 15:21 adev Accepted_Events_Run_2064
  • 33375 02/10/00 14:43 adev Accepted_Events_Run_2063
  • 33372 02/10/00 14:37 adev Accepted_Events_Run_2062
  • 33369 02/10/00 14:30 adev Accepted_Events_Run_2061
  • 33366 02/10/00 14:17 adev Accepted_Events_Run_2060
  • 33364 02/10/00 14:01 Saha Beam striptool
  • 33362 02/10/00 14:00 Saha Magnet striptools
  • 33361 02/10/00 13:00 adev Accepted_Events_Run_2059
  • 33359 02/10/00 12:46 S.Malov magnets status
  • 33357 02/10/00 12:20 Saha Harp Scan - No Raster
  • 33356 02/10/00 12:15 A. Saha Event sum for #35
  • 33355 02/10/00 12:10 adev Accepted_Events_Run_2058
  • 33352 02/10/00 12:05 adev Accepted_Events_Run_2057
  • 33349 02/10/00 11:37 adev Accepted_Events_Run_2056
  • 33346 02/10/00 10:56 adev Accepted_Events_Run_2054
  • 33344 02/10/00 10:43 Saha Beam conditions after PSS down
  • 33342 02/10/00 10:29 S.Malov spot size
  • 33341 02/10/00 10:07 Saha Hall in restricted Access
  • 33340 02/10/00 08:19 M. Rvachev E_miss plot for kin 35
  • 33339 02/10/00 08:10 M. Rvachev Shift summary
  • 33337 02/10/00 07:48 M. Rvachev Beam stripcharts
  • 33336 02/10/00 07:36 M. Rvachev Current history
  • 33334 02/10/00 07:12 adev Accepted_Events_Run_2053
  • 33333 02/10/00 06:43 adev Accepted_Events_Run_2052
  • 33330 02/10/00 05:48 adev Accepted_Events_Run_2051
  • 33327 02/10/00 05:04 Eugene Raster size
  • 33325 02/10/00 04:51 adev Accepted_Events_Run_2050
  • 33322 02/10/00 04:49 adev Accepted_Events_Run_2049
  • 33318 02/10/00 04:45 adev Accepted_Events_Run_2048
  • 33317 02/10/00 04:42 M. Rvachev Changed to kin 35
  • 33314 02/10/00 04:16 adev Accepted_Events_Run_2047
  • 33313 02/10/00 04:08 gayou stripcharts
  • 33311 02/10/00 04:06 gayou Stripcharts
  • 33310 02/10/00 03:56 Eugene The raster is repared
  • 33309 02/10/00 03:13 M. Rvachev Raster problems
  • 33308 02/10/00 03:11 Nilanga Event sum for # 34
  • 33306 02/10/00 02:25 adev Accepted_Events_Run_2045
  • 33304 02/10/00 02:01 adev Accepted_Events_Run_2044
  • 33302 02/10/00 01:37 adev Accepted_Events_Run_2043
  • 33299 02/10/00 00:46 Nilanga Good events for #33
  • 33298 02/10/00 00:40 adev Accepted_Events_Run_2042
  • 33295 02/10/00 00:17 Nilanga Cuts for # 35 and #36
  • 33294 02/10/00 00:11 LeRose Shift Summary
  • Wednesday

  • 33292 02/09/00 23:42 adev Accepted_Events_Run_2041
  • 33289 02/09/00 23:20 LeRose Beam Strip Chart
  • 33288 02/09/00 23:19 LeRose OTR
  • 33287 02/09/00 22:49 LeRose Beam Current
  • 33285 02/09/00 22:40 adev Accepted_Events_Run_2040
  • 33282 02/09/00 22:16 LeRose Kinematics 34
  • 33281 02/09/00 21:55 LeRose End Kinematics 33
  • 33280 02/09/00 21:53 adev Accepted_Events_Run_2039
  • 33277 02/09/00 21:37 adev Accepted_Events_Run_2038
  • 33273 02/09/00 21:26 adev Accepted_Events_Run_2037
  • 33271 02/09/00 21:06 adev Accepted_Events_Run_2036
  • 33267 02/09/00 20:31 adev Accepted_Events_Run_2035
  • 33265 02/09/00 20:18 adev Accepted_Events_Run_2034
  • 33263 02/09/00 20:11 LeRose Beam Strip Chart
  • 33261 02/09/00 19:26 adev Accepted_Events_Run_2033
  • 33257 02/09/00 19:12 adev Accepted_Events_Run_2032
  • 33255 02/09/00 19:05 adev Accepted_Events_Run_2031
  • 33253 02/09/00 19:01 adev Accepted_Events_Run_2030
  • 33249 02/09/00 18:28 Nilanga Elastic scattering at 5 fm^-2
  • 33248 02/09/00 18:12 Saha OTR in/out - Hall A harp scans
  • 33247 02/09/00 17:30 LeRose Electron Arm Reset
  • 33246 02/09/00 16:04 adev Accepted_Events_Run_2029
  • 33244 02/09/00 16:02 Serge Kox Shift Summary 08-16h 2/9/00
  • 33241 02/09/00 15:54 adev Accepted_Events_Run_2028
  • 33238 02/09/00 15:49 Higinbotham Elastic Scattering at 5.0 fm-2
  • 33237 02/09/00 15:43 adev Accepted_Events_Run_2027
  • 33236 02/09/00 15:43 Hansen Summary for runs 1973-2021
  • 33233 02/09/00 15:30 adev Accepted_Events_Run_2026
  • 33229 02/09/00 15:11 adev Accepted_Events_Run_2025
  • 33227 02/09/00 15:06 adev Accepted_Events_Run_2024
  • 33224 02/09/00 15:05 adev Accepted_Events_Run_2023
  • 33220 02/09/00 15:02 Adaq Spectrometer setting
  • 33219 02/09/00 13:07 adev Accepted_Events_Run_2022
  • 33216 02/09/00 12:50 adev Accepted_Events_Run_2021
  • 33214 02/09/00 12:45 Liatard magnet strip charts
  • 33213 02/09/00 12:39 Liatard beam strip chart
  • 33212 02/09/00 12:13 adev Accepted_Events_Run_2020
  • 33208 02/09/00 10:25 Serge Kox (French) Harp scan Voutier/Kox
  • 33206 02/09/00 10:19 adev Accepted_Events_Run_2019
  • 33204 02/09/00 10:14 adev Accepted_Events_Run_2018
  • 33201 02/09/00 09:52 adev Accepted_Events_Run_2017
  • 33198 02/09/00 08:43 adev Accepted_Events_Run_2016
  • 33196 02/09/00 08:02 Eric Voutier Shift Summary
  • 33194 02/09/00 07:26 adev Accepted_Events_Run_2015
  • 33191 02/09/00 06:00 adev Accepted_Events_Run_2014
  • 33188 02/09/00 05:20 Eric Voutier Magnet strip chart
  • 33187 02/09/00 05:03 Eric Voutier Beam strip chart
  • 33186 02/09/00 04:59 Eric Voutier Magnet settings
  • 33184 02/09/00 03:44 kuss run #2012
  • 33183 02/09/00 01:26 kuss quadrupole cycling
  • 33182 02/09/00 00:53 adev Accepted_Events_Run_2013
  • 33180 02/09/00 00:47 kuss fan speed
  • 33179 02/09/00 00:39 adev Accepted_Events_Run_2012
  • 33176 02/09/00 00:07 Hansen Shift summary
  • Tuesday

  • 33174 02/08/00 23:24 adev Accepted_Events_Run_2011
  • 33171 02/08/00 22:00 punjabi Fan on Loop 1
  • 33170 02/08/00 21:49 adev Accepted_Events_Run_2010
  • 33167 02/08/00 21:43 punjabi Fan on Loop1
  • 33166 02/08/00 21:12 Hansen E89-044 strip chart
  • 33165 02/08/00 21:11 Hansen Magnet strip charts
  • 33164 02/08/00 20:16 Hansen Beam energy spread, OTR out
  • 33163 02/08/00 20:14 adev Accepted_Events_Run_2009
  • 33158 02/08/00 18:29 adev Accepted_Events_Run_2008
  • 33156 02/08/00 18:08 adev Accepted_Events_Run_2007
  • 33153 02/08/00 17:52 adev Accepted_Events_Run_2006
  • 33151 02/08/00 17:44 Tieulent Raster size with harp
  • 33150 02/08/00 17:44 Hansen Raster spot
  • 33148 02/08/00 17:40 Tieulent Harp Scan
  • 33146 02/08/00 17:32 adev Accepted_Events_Run_2005
  • 33144 02/08/00 17:31 adev Accepted_Events_Run_2004
  • 33140 02/08/00 16:28 adev Accepted_Events_Run_2003
  • 33138 02/08/00 16:08 Saha Bdl Measurement
  • 33136 02/08/00 12:54 Saha Beam Update
  • 33134 02/08/00 11:48 adev Accepted_Events_Run_2002
  • 33133 02/08/00 11:44 kuss target status
  • 33131 02/08/00 11:25 Saha E89044 - Beam Related Checks and Record during Shift
  • 33129 02/08/00 10:22 adev Accepted_Events_Run_2001
  • 33128 02/08/00 10:21 R. Michaels pedrun config works again
  • 33126 02/08/00 10:20 fatiha Target chart
  • 33124 02/08/00 10:14 adev Accepted_Events_Run_1999
  • 33122 02/08/00 10:06 adev Accepted_Events_Run_1997
  • 33120 02/08/00 10:03 R. Michaels probably solution to ER1 woes
  • 33119 02/08/00 10:02 fatiha Target problem
  • 33116 02/08/00 08:36 kuss Emiss for #1856
  • 33115 02/08/00 08:19 kuss Emiss for #1856 (kin #21)
  • 33113 02/08/00 08:05 M. Rvachev Shift summary
  • 33112 02/08/00 07:39 Liatard Stripcharts
  • 33111 02/08/00 07:38 M. Rvachev 2bbu rate estimate
  • 33110 02/08/00 06:52 M. Rvachev 2bbu rate estimate
  • 33108 02/08/00 05:35 adev Accepted_Events_Run_1996
  • 33107 02/08/00 05:25 Liatard OTR
  • 33105 02/08/00 05:19 adev Accepted_Events_Run_1995
  • 33103 02/08/00 04:08 adev Accepted_Events_Run_1994
  • 33101 02/08/00 03:57 frullani OTR screen
  • 33099 02/08/00 03:54 adev Accepted_Events_Run_1993
  • 33097 02/08/00 03:13 Liatard beam spot
  • 33094 02/08/00 02:52 adev Accepted_Events_Run_1992
  • 33093 02/08/00 02:40 M. Rvachev OTR screen
  • 33091 02/08/00 02:39 adev Accepted_Events_Run_1991
  • 33087 02/08/00 01:32 adev Accepted_Events_Run_1990
  • 33086 02/08/00 01:22 M. Rvachev OTR screen
  • 33084 02/08/00 01:17 adev Accepted_Events_Run_1989
  • 33082 02/08/00 01:05 Liatard Beam strip chart
  • 33081 02/08/00 01:03 Liatard Beam strip chart
  • 33080 02/08/00 00:22 Eric Voutier Raster size
  • 33078 02/08/00 00:19 Eric Voutier Shift summary
  • 33076 02/08/00 00:06 adev Accepted_Events_Run_1988
  • Monday

  • 33073 02/07/00 23:32 Hansen Spot++
  • 33072 02/07/00 23:30 adev Accepted_Events_Run_1987
  • 33069 02/07/00 23:26 adev Accepted_Events_Run_1986
  • 33066 02/07/00 23:22 adev Accepted_Events_Run_1985
  • 33063 02/07/00 23:19 adev Accepted_Events_Run_1984
  • 33060 02/07/00 23:12 adev Accepted_Events_Run_1983
  • 33057 02/07/00 23:07 adev Accepted_Events_Run_1982
  • 33055 02/07/00 22:34 Hansen Runs 1980-1982 are junk
  • 33053 02/07/00 21:42 adev Accepted_Events_Run_1981
  • 33052 02/07/00 21:02 R. Michaels scaler oddities, rate observation
  • 33048 02/07/00 16:37 adev Accepted_Events_Run_1979
  • 33045 02/07/00 16:11 Calarco Shift Summary
  • 33044 02/07/00 16:10 Calarco Run record for kin #25
  • 33043 02/07/00 15:41 jpchen Target is out of beam
  • 33041 02/07/00 14:06 jpchen target ioc reboot
  • 33040 02/07/00 13:47 adev Accepted_Events_Run_1977
  • 33037 02/07/00 13:10 tieulent OTR Check
  • 33036 02/07/00 13:09 adev Accepted_Events_Run_1976
  • 33033 02/07/00 12:06 adev Accepted_Events_Run_1975
  • 33032 02/07/00 12:06 tieulent otr check
  • 33030 02/07/00 11:36 Tieulent Beam parameters (rescale)
  • 33029 02/07/00 11:36 R. Michaels More about BPM (satellite peaks, resol)
  • 33028 02/07/00 11:05 tieulent beam parameters
  • 33026 02/07/00 10:58 Calarco OTR check
  • 33025 02/07/00 10:43 adev Accepted_Events_Run_1974
  • 33024 02/07/00 10:43 R. Michaels T4 rate
  • 33022 02/07/00 10:17 R. Michaels BPM resolution, ADC problems
  • 33020 02/07/00 09:38 adev Accepted_Events_Run_1973
  • 33018 02/07/00 09:00 Tieulent Spot
  • 33016 02/07/00 08:55 tieulent Spec setting kin 25
  • 33015 02/07/00 08:53 Calarco OTR
  • 33014 02/07/00 08:50 markowitz OTR
  • 33013 02/07/00 08:48 adev Accepted_Events_Run_1972
  • 33010 02/07/00 08:47 adev Accepted_Events_Run_1971
  • 33006 02/07/00 08:14 markowitz shift summary
  • 33005 02/07/00 08:02 adev Accepted_Events_Run_1970
  • 33002 02/07/00 07:59 adev Accepted_Events_Run_1969
  • 32999 02/07/00 07:46 adaq e89044 beam
  • 32998 02/07/00 07:43 adev Accepted_Events_Run_1968
  • 32994 02/07/00 06:52 adev Accepted_Events_Run_1967
  • 32991 02/07/00 05:31 adev Accepted_Events_Run_1966
  • 32989 02/07/00 04:57 markowitz magnet
  • 32987 02/07/00 04:40 adev Accepted_Events_Run_1965
  • 32984 02/07/00 04:30 adev Accepted_Events_Run_1964
  • 32981 02/07/00 04:23 adev Accepted_Events_Run_1963
  • 32978 02/07/00 03:59 adev Accepted_Events_Run_1962
  • 32976 02/07/00 02:47 frullani kinematical point 19
  • 32975 02/07/00 02:38 frullani beam spot
  • 32973 02/07/00 02:05 adev Accepted_Events_Run_1961
  • 32970 02/07/00 01:32 fatiha Elastic rate
  • 32968 02/07/00 00:32 Higinbotham Shift Summary
  • 32967 02/07/00 00:26 adev Accepted_Events_Run_1960
  • 32965 02/07/00 00:12 adev Accepted_Events_Run_1959
  • Sunday

  • 32962 02/06/00 23:57 adev Accepted_Events_Run_1957
  • 32961 02/06/00 23:54 adev Accepted_Events_Run_1958
  • 32955 02/06/00 23:37 adev Accepted_Events_Run_1956
  • 32953 02/06/00 23:34 adev Accepted_Events_Run_1955
  • 32950 02/06/00 23:29 adev Accepted_Events_Run_1954
  • 32947 02/06/00 23:23 adev Accepted_Events_Run_1952
  • 32946 02/06/00 23:22 adev Accepted_Events_Run_1952
  • 32944 02/06/00 23:11 adev Accepted_Events_Run_1953
  • 32938 02/06/00 22:56 adev Accepted_Events_Run_1951
  • 32936 02/06/00 22:46 adev Accepted_Events_Run_1950
  • 32933 02/06/00 22:39 adev Accepted_Events_Run_1949
  • 32929 02/06/00 22:23 adev Accepted_Events_Run_1948
  • 32928 02/06/00 22:22 mitchell ep-energy
  • 32926 02/06/00 21:03 Higinbotham Header Files
  • 32925 02/06/00 20:51 Saha Arc Energy Measurement, Feb 6,2000
  • 32924 02/06/00 18:06 Liatard Beam Optics Stripchart
  • 32923 02/06/00 18:02 Higinbotham Beam Strip Charts
  • 32921 02/06/00 17:23 adev Accepted_Events_Run_1947
  • 32919 02/06/00 17:23 adev Accepted_Events_Run_1946
  • 32915 02/06/00 16:44 R. Michaels DAQ problems, possible solutions
  • 32914 02/06/00 16:41 adev Accepted_Events_Run_1945
  • 32912 02/06/00 16:30 adev Accepted_Events_Run_1944
  • 32907 02/06/00 16:21 adev Accepted_Events_Run_1943
  • 32906 02/06/00 16:20 adev Accepted_Events_Run_1942
  • 32901 02/06/00 16:00 Calarco Shift summary
  • 32900 02/06/00 15:58 adev Accepted_Events_Run_1940
  • 32898 02/06/00 15:52 adev Accepted_Events_Run_1939
  • 32895 02/06/00 13:56 Saha Arc-Energy Measurement, Feb 3, 2000
  • 32891 02/06/00 12:43 adev Accepted_Events_Run_1936
  • 32889 02/06/00 10:41 adev Accepted_Events_Run_1935
  • 32887 02/06/00 10:34 Nilanga 12C runs at 2.9 GeV
  • 32885 02/06/00 09:12 adev Accepted_Events_Run_1934
  • 32881 02/06/00 08:28 adev Accepted_Events_Run_1933
  • 32880 02/06/00 08:06 markowitz shift summary
  • 32878 02/06/00 07:34 adev Accepted_Events_Run_1932
  • 32875 02/06/00 06:53 adev Accepted_Events_Run_1931
  • 32870 02/06/00 06:32 adev Accepted_Events_Run_1929
  • 32869 02/06/00 06:32 adev Accepted_Events_Run_1930
  • 32865 02/06/00 05:29 adev Accepted_Events_Run_1928
  • 32862 02/06/00 05:19 adev Accepted_Events_Run_1927
  • 32858 02/06/00 05:08 adev Accepted_Events_Run_1925
  • 32855 02/06/00 04:59 adev Accepted_Events_Run_1924
  • 32852 02/06/00 04:46 adev Accepted_Events_Run_1923
  • 32850 02/06/00 04:33 adev Accepted_Events_Run_1922
  • 32846 02/06/00 04:13 adev Accepted_Events_Run_1920
  • 32843 02/06/00 03:56 adev Accepted_Events_Run_1919
  • 32840 02/06/00 03:42 markowitz e89044 beam
  • 32838 02/06/00 03:34 adev Accepted_Events_Run_1918
  • 32835 02/06/00 03:32 adev Accepted_Events_Run_1917
  • 32833 02/06/00 03:28 adev Accepted_Events_Run_1915
  • 32828 02/06/00 03:04 adev Accepted_Events_Run_1914
  • 32826 02/06/00 03:03 adev Accepted_Events_Run_1913
  • 32823 02/06/00 02:48 adev Accepted_Events_Run_1912
  • 32819 02/06/00 02:15 adev Accepted_Events_Run_1910
  • 32816 02/06/00 01:59 Tieulent Spot ++
  • 32814 02/06/00 01:50 adev Accepted_Events_Run_1909
  • 32811 02/06/00 01:06 adev Accepted_Events_Run_1907
  • 32810 02/06/00 01:06 adev Accepted_Events_Run_1907
  • 32809 02/06/00 01:05 adev Accepted_Events_Run_1908
  • 32804 02/06/00 00:32 adev Accepted_Events_Run_1906
  • 32801 02/06/00 00:27 markowitz e89044 beam
  • 32800 02/06/00 00:26 markowitz beam and magnets
  • 32799 02/06/00 00:21 markowitz otr+accelerator
  • 32795 02/06/00 00:03 Adaq Shift Summary
  • 32794 02/06/00 00:03 adev Accepted_Events_Run_1904
  • 32791 02/06/00 00:01 adev Accepted_Events_Run_1903
  • Saturday

  • 32787 02/05/00 23:59 adev Accepted_Events_Run_1902
  • 32785 02/05/00 23:58 adev Accepted_Events_Run_1901
  • 32783 02/05/00 23:57 fatiha Total number of event
  • 32781 02/05/00 23:52 adev Accepted_Events_Run_1899
  • 32778 02/05/00 23:32 adev Accepted_Events_Run_1898
  • 32775 02/05/00 23:10 adev Accepted_Events_Run_1897
  • 32771 02/05/00 23:07 adev Accepted_Events_Run_1895
  • 32770 02/05/00 23:06 adev Accepted_Events_Run_1896
  • 32766 02/05/00 23:05 adev Accepted_Events_Run_1893
  • 32765 02/05/00 23:03 adev Accepted_Events_Run_1894
  • 32760 02/05/00 22:46 adev Accepted_Events_Run_1892
  • 32757 02/05/00 22:16 adev Accepted_Events_Run_1890
  • 32753 02/05/00 21:58 adev Accepted_Events_Run_1889
  • 32751 02/05/00 21:45 Higinbotham OTR
  • 32750 02/05/00 21:44 adev Accepted_Events_Run_1888
  • 32746 02/05/00 21:22 adev Accepted_Events_Run_1887
  • 32743 02/05/00 21:12 Higinbotham Magnet Settings
  • 32742 02/05/00 21:11 Higinbotham strip charts
  • 32741 02/05/00 21:04 adev Accepted_Events_Run_1886
  • 32738 02/05/00 20:45 adev Accepted_Events_Run_1885
  • 32735 02/05/00 20:38 adev Accepted_Events_Run_1884
  • 32732 02/05/00 20:19 adev Accepted_Events_Run_1883
  • 32729 02/05/00 20:00 adev Accepted_Events_Run_1882
  • 32726 02/05/00 19:45 Fatiha Electron pion separation
  • 32725 02/05/00 19:35 adev Accepted_Events_Run_1881
  • 32721 02/05/00 19:34 adev Accepted_Events_Run_1880
  • 32718 02/05/00 19:32 adev Accepted_Events_Run_1879
  • 32715 02/05/00 19:11 adev Accepted_Events_Run_1878
  • 32711 02/05/00 18:59 adev Accepted_Events_Run_1877
  • 32709 02/05/00 18:33 fatiha Reboot Coda
  • 32708 02/05/00 18:30 adev Accepted_Events_Run_1876
  • 32705 02/05/00 18:06 fatiha Spot size
  • 32702 02/05/00 17:59 adev Accepted_Events_Run_1874
  • 32701 02/05/00 17:54 fatiha Target chart 1874
  • 32698 02/05/00 17:37 adev Accepted_Events_Run_1873
  • 32697 02/05/00 17:16 adev Accepted_Events_Run_1872
  • 32693 02/05/00 17:02 adev Accepted_Events_Run_1871
  • 32689 02/05/00 16:38 adev Accepted_Events_Run_1870
  • 32687 02/05/00 16:12 Calarco Shift Summary
  • 32685 02/05/00 15:30 kuss rates for kin #21
  • 32684 02/05/00 15:18 kuss resolution of 1856
  • 32683 02/05/00 14:46 Calarco beam history
  • 32681 02/05/00 14:10 adev Accepted_Events_Run_1869
  • 32679 02/05/00 14:07 adev Accepted_Events_Run_1868
  • 32676 02/05/00 13:03 kuss new trigsetup loaded
  • 32675 02/05/00 13:02 adev Accepted_Events_Run_1866
  • 32674 02/05/00 12:59 kuss hac gui
  • 32672 02/05/00 12:51 adev Accepted_Events_Run_1865
  • 32667 02/05/00 11:49 adev Accepted_Events_Run_1864
  • 32666 02/05/00 11:48 kuss target tests
  • 32663 02/05/00 11:15 adev Accepted_Events_Run_1863
  • 32660 02/05/00 10:59 adev Accepted_Events_Run_1862
  • 32657 02/05/00 10:40 adev Accepted_Events_Run_1861
  • 32654 02/05/00 10:15 adev Accepted_Events_Run_1860
  • 32652 02/05/00 09:49 kuss target sleeps
  • 32651 02/05/00 09:47 kuss end of kin #21
  • 32650 02/05/00 09:39 adev Accepted_Events_Run_1859
  • 32648 02/05/00 08:17 Eric Voutier Shift summary
  • 32646 02/05/00 08:06 adev Accepted_Events_Run_1858
  • 32643 02/05/00 07:59 Adaq otr cabling
  • 32642 02/05/00 06:37 adev Accepted_Events_Run_1857
  • 32639 02/05/00 05:45 Eric Voutier Beam strip chart
  • 32638 02/05/00 05:00 adev Accepted_Events_Run_1856
  • 32633 02/05/00 02:43 Adaq daq problems
  • 32631 02/05/00 02:38 adev Accepted_Events_Run_1852
  • 32628 02/05/00 01:05 adev Accepted_Events_Run_1851
  • 32625 02/05/00 00:34 markowitz otr
  • 32624 02/05/00 00:33 markowitz accelerator bpm screens
  • 32623 02/05/00 00:31 markowitz magnet screen
  • 32622 02/05/00 00:29 markowitz beam parameters and current
  • 32619 02/05/00 00:04 Tieulent Shift summary
  • Friday

  • 32616 02/04/00 23:17 adev Accepted_Events_Run_1850
  • 32614 02/04/00 22:46 roblin ep cerenkov fixed
  • 32613 02/04/00 21:17 Tieulent Magnet stripchart
  • 32612 02/04/00 20:45 Tieulent Beam Parameters
  • 32610 02/04/00 19:10 adev Accepted_Events_Run_1848
  • 32608 02/04/00 19:06 Tieulent Problem in the injector
  • 32607 02/04/00 18:36 M. Rvachev Missing End of Run comments
  • 32603 02/04/00 18:18 Tieulent Spot
  • 32601 02/04/00 18:17 adev Accepted_Events_Run_1847
  • 32587 02/04/00 17:52 adev Accepted_Events_Run_1846
  • 32586 02/04/00 17:52 Tieulent Spot ++
  • 32583 02/04/00 17:37 Tieulent Change electron setting
  • 32579 02/04/00 16:22 Calarco beam parameters
  • 32578 02/04/00 16:13 Saha Shift Summary
  • 32576 02/04/00 16:00 adev Accepted_Events_Run_1844
  • 32573 02/04/00 15:52 adev Accepted_Events_Run_1843
  • 32572 02/04/00 15:40 Calarco spot
  • 32570 02/04/00 15:36 R. Michaels predicted deadtime with high prescale
  • 32567 02/04/00 15:13 adev Accepted_Events_Run_1842
  • 32563 02/04/00 13:56 adev Accepted_Events_Run_1841
  • 32561 02/04/00 13:44 Calarco dhist
  • 32560 02/04/00 13:35 Saha OTR Screen and Maxvideo
  • 32559 02/04/00 13:33 Saha Beam Parameers
  • 32558 02/04/00 13:32 adev Accepted_Events_Run_1840
  • 32557 02/04/00 13:31 Saha HAC Screen
  • 32554 02/04/00 13:12 adev Accepted_Events_Run_1839
  • 32551 02/04/00 12:46 adev Accepted_Events_Run_1838
  • 32548 02/04/00 12:30 R. Michaels DAQ problem (ER1, ROC14)
  • 32547 02/04/00 12:26 doug Elastic Rate Check
  • 32546 02/04/00 12:23 adev Accepted_Events_Run_1837
  • 32542 02/04/00 12:10 adev Accepted_Events_Run_1836
  • 32540 02/04/00 11:59 adev Accepted_Events_Run_1835
  • 32537 02/04/00 11:52 adev Accepted_Events_Run_1834
  • 32534 02/04/00 11:50 adev Accepted_Events_Run_1833
  • 32529 02/04/00 11:36 adev Accepted_Events_Run_1831
  • 32525 02/04/00 11:26 adev Accepted_Events_Run_1830
  • 32522 02/04/00 11:18 adev Accepted_Events_Run_1829
  • 32520 02/04/00 11:09 adev Accepted_Events_Run_1828
  • 32517 02/04/00 10:59 adev Accepted_Events_Run_1827
  • 32515 02/04/00 10:53 adev Accepted_Events_Run_1826
  • 32512 02/04/00 10:41 adev Accepted_Events_Run_1825
  • 32509 02/04/00 10:27 adev Accepted_Events_Run_1824
  • 32506 02/04/00 10:20 adev Accepted_Events_Run_1821
  • 32505 02/04/00 10:20 adev Accepted_Events_Run_1821
  • 32503 02/04/00 10:03 adev Accepted_Events_Run_1823
  • 32502 02/04/00 10:03 R. Michaels CODA problems
  • 32499 02/04/00 10:00 adev Accepted_Events_Run_1822
  • 32492 02/04/00 09:37 adev Accepted_Events_Run_1819
  • 32488 02/04/00 09:33 adev Accepted_Events_Run_1817
  • 32486 02/04/00 09:28 adev Accepted_Events_Run_1816
  • 32483 02/04/00 09:27 adev Accepted_Events_Run_1815
  • 32479 02/04/00 08:50 Saha OTR Screen, Beam Current
  • 32478 02/04/00 08:49 adev Accepted_Events_Run_1814
  • 32477 02/04/00 08:48 Saha Beam Parameters
  • 32471 02/04/00 08:14 adev Accepted_Events_Run_1810
  • 32468 02/04/00 08:10 adev Accepted_Events_Run_1809
  • 32466 02/04/00 08:07 Adev more for bob/daq troubles
  • 32464 02/04/00 08:04 fatiha Elastic rate for run number 1792
  • 32462 02/04/00 07:46 adev Accepted_Events_Run_1808
  • 32459 02/04/00 07:33 adev Accepted_Events_Run_1807
  • 32456 02/04/00 07:32 adev Accepted_Events_Run_1806
  • 32454 02/04/00 07:32 Adev for rom, paste of screen messages from hung daq
  • 32453 02/04/00 07:31 adev Accepted_Events_Run_1805
  • 32450 02/04/00 07:16 adev Accepted_Events_Run_1804
  • 32446 02/04/00 06:57 adev Accepted_Events_Run_1803
  • 32444 02/04/00 06:55 adev Accepted_Events_Run_1802
  • 32441 02/04/00 06:41 adev Accepted_Events_Run_1801
  • 32438 02/04/00 06:26 adev Accepted_Events_Run_1800
  • 32435 02/04/00 06:13 markowitz daq troubles, junk runs
  • 32432 02/04/00 05:48 adev Accepted_Events_Run_1798
  • 32429 02/04/00 05:45 adev Accepted_Events_Run_1796
  • 32428 02/04/00 05:45 adev Accepted_Events_Run_1797
  • 32425 02/04/00 05:42 markowitz daq troubles
  • 32423 02/04/00 05:38 markowitz otr
  • 32422 02/04/00 05:36 adev Accepted_Events_Run_1795
  • 32419 02/04/00 05:28 fatiha Target chart 1795
  • 32417 02/04/00 05:18 adev Accepted_Events_Run_1794
  • 32415 02/04/00 05:08 fatiha Target chart 1794
  • 32414 02/04/00 05:05 adev Accepted_Events_Run_1793
  • 32411 02/04/00 04:55 fatiha Traget chart 1793
  • 32408 02/04/00 04:47 adev Accepted_Events_Run_1792
  • 32407 02/04/00 04:35 fatiha Target chart 1792
  • 32405 02/04/00 04:31 adev Accepted_Events_Run_1791
  • 32403 02/04/00 04:09 fatiha Target chart 1791
  • 32401 02/04/00 04:04 markowitz finished I-scan, preparing elastic
  • 32400 02/04/00 03:59 adev Accepted_Events_Run_1790
  • 32398 02/04/00 03:53 fatiha Target chart 1790
  • 32396 02/04/00 03:41 adev Accepted_Events_Run_1789
  • 32395 02/04/00 03:40 fatiha Target chart 1789
  • 32392 02/04/00 03:34 adev Accepted_Events_Run_1788
  • 32389 02/04/00 03:19 adev Accepted_Events_Run_1787
  • 32386 02/04/00 03:10 fatiha Target chart 1787
  • 32385 02/04/00 03:08 R. Michaels Av deadtime inconsistent
  • 32384 02/04/00 03:03 adev Accepted_Events_Run_1786
  • 32381 02/04/00 02:43 Adaq current monitor and magnets screen
  • 32380 02/04/00 02:42 adev Accepted_Events_Run_1785
  • 32378 02/04/00 02:40 fatiha Target chart 1785
  • 32377 02/04/00 02:37 adaq general accelerator parameters
  • 32375 02/04/00 02:35 adaq otr
  • 32374 02/04/00 02:34 Adaq he e89044 beam current
  • 32373 02/04/00 02:25 adev Accepted_Events_Run_1784
  • 32371 02/04/00 02:20 fatiha Target chart 1784
  • 32369 02/04/00 02:18 adev Accepted_Events_Run_1783
  • 32366 02/04/00 02:11 adev Accepted_Events_Run_1782
  • 32363 02/04/00 02:06 adev Accepted_Events_Run_1781
  • 32360 02/04/00 01:57 adev Accepted_Events_Run_1780
  • 32359 02/04/00 01:57 fatiha Target chart 1780
  • 32356 02/04/00 01:39 fatiha Target chart 1779
  • 32355 02/04/00 01:34 adev Accepted_Events_Run_1779
  • 32352 02/04/00 01:19 adev Accepted_Events_Run_1778
  • 32349 02/04/00 01:02 adev Accepted_Events_Run_1777
  • 32347 02/04/00 00:56 fatiha Target chart 1777
  • 32345 02/04/00 00:48 adev Accepted_Events_Run_1776
  • 32342 02/04/00 00:35 adev Accepted_Events_Run_1775
  • 32340 02/04/00 00:32 fatiha Target chart 1775
  • 32338 02/04/00 00:22 adev Accepted_Events_Run_1774
  • 32336 02/04/00 00:16 fatiha target chart 1774
  • 32334 02/04/00 00:05 adev Accepted_Events_Run_1773
  • Thursday

  • 32331 02/03/00 23:56 adev Accepted_Events_Run_1772
  • 32329 02/03/00 23:54 fatiha target chart
  • 32325 02/03/00 23:43 Eric Voutier Current Scan
  • 32324 02/03/00 23:35 adev Accepted_Events_Run_1771
  • 32321 02/03/00 23:19 adev Accepted_Events_Run_1770
  • 32318 02/03/00 23:05 adev Accepted_Events_Run_1769
  • 32315 02/03/00 22:54 Jack Resets
  • 32313 02/03/00 22:48 adev Accepted_Events_Run_1768
  • 32311 02/03/00 22:38 adev Accepted_Events_Run_1767
  • 32308 02/03/00 22:30 adev Accepted_Events_Run_1766
  • 32305 02/03/00 22:16 adev Accepted_Events_Run_1765
  • 32302 02/03/00 22:08 adev Accepted_Events_Run_1764
  • 32299 02/03/00 22:00 adev Accepted_Events_Run_1763
  • 32297 02/03/00 21:57 punjabi spot++
  • 32295 02/03/00 21:50 adev Accepted_Events_Run_1762
  • 32291 02/03/00 21:41 adev Accepted_Events_Run_1761
  • 32289 02/03/00 21:35 adev Accepted_Events_Run_1760
  • 32286 02/03/00 21:25 adev Accepted_Events_Run_1759
  • 32284 02/03/00 20:38 Eric Voutier OTR strip chart
  • 32282 02/03/00 20:34 Eric Voutier Beam status
  • 32281 02/03/00 20:21 adev Accepted_Events_Run_1758
  • 32278 02/03/00 19:08 adev Accepted_Events_Run_1757
  • 32277 02/03/00 19:08 kuss spot for 1757
  • 32275 02/03/00 19:05 adev Accepted_Events_Run_1756
  • 32274 02/03/00 19:05 kuss spot for 1756
  • 32271 02/03/00 18:23 adev Accepted_Events_Run_1755
  • 32269 02/03/00 18:22 kuss spot for 1755
  • 32267 02/03/00 18:15 adev Accepted_Events_Run_1754
  • 32264 02/03/00 18:13 adev Accepted_Events_Run_1753
  • 32261 02/03/00 17:57 fatiha Target chart
  • 32259 02/03/00 17:47 adev Accepted_Events_Run_1752
  • 32258 02/03/00 17:46 kuss spot for 1752
  • 32255 02/03/00 17:43 adev Accepted_Events_Run_1751
  • 32254 02/03/00 17:42 kuss spot for 1751
  • 32252 02/03/00 17:37 kuss spot for 1750 #2
  • 32251 02/03/00 17:36 kuss spot for 1750
  • 32249 02/03/00 17:35 adev Accepted_Events_Run_1750
  • 32247 02/03/00 17:22 kuss spot for 1749
  • 32245 02/03/00 17:20 adev Accepted_Events_Run_1749
  • 32243 02/03/00 17:16 kuss spot for 1746
  • 32242 02/03/00 17:15 kuss spot for 1747
  • 32240 02/03/00 17:11 adev Accepted_Events_Run_1748
  • 32239 02/03/00 17:11 kuss spot for 1748
  • 32236 02/03/00 16:59 adev Accepted_Events_Run_1747
  • 32233 02/03/00 16:54 adev Accepted_Events_Run_1746
  • 32231 02/03/00 16:33 adev Accepted_Events_Run_1745
  • 32226 02/03/00 15:54 adev Accepted_Events_Run_1744
  • 32224 02/03/00 15:35 adev Accepted_Events_Run_1743
  • 32221 02/03/00 15:34 adev Accepted_Events_Run_1742
  • 32217 02/03/00 15:19 adev Accepted_Events_Run_1741
  • 32215 02/03/00 15:17 adev Accepted_Events_Run_1740
  • 32212 02/03/00 15:14 R. Michaels dropped ROC13
  • 32211 02/03/00 15:07 adev Accepted_Events_Run_1738
  • 32208 02/03/00 14:49 adev Accepted_Events_Run_1735
  • 32204 02/03/00 14:32 adev Accepted_Events_Run_1733
  • 32200 02/03/00 14:07 adev Accepted_Events_Run_1732
  • 32197 02/03/00 13:53 adev Accepted_Events_Run_1731
  • 32194 02/03/00 13:07 adev Accepted_Events_Run_1730
  • 32192 02/03/00 11:34 Margaziotis shift status
  • 32190 02/03/00 10:07 adev Accepted_Events_Run_1729
  • 32188 02/03/00 08:18 roblin EP measurement
  • 32187 02/03/00 08:14 kuss trigger setup HRSH 2.903GeV
  • 32186 02/03/00 07:51 Saha Beam Parameters
  • 32185 02/03/00 05:51 Tieulent Starting ep measurement
  • 32184 02/03/00 04:14 mitchell ep status
  • 32183 02/03/00 04:09 mitchell Compton - summary
  • 32182 02/03/00 02:48 Tieulent Compton Started
  • Wednesday

  • 32181 02/02/00 22:52 Margaziotis Continued beam problems
  • 32180 02/02/00 22:05 Margaziotis Compton start
  • 32179 02/02/00 22:00 Margaziotis No CW beam
  • 32178 02/02/00 21:06 Margaziotis Compton
  • 32177 02/02/00 20:04 Jack Resets
  • 32176 02/02/00 19:37 Margaziotis Beam update
  • 32175 02/02/00 19:14 Margaziotis Beam status
  • 32174 02/02/00 16:33 Margaziotis Hall lockup
  • 32171 02/02/00 14:53 adev Accepted_Events_Run_1727
  • 32169 02/02/00 14:12 adev Accepted_Events_Run_1726
  • Tuesday

  • 32164 02/01/00 16:40 adev Accepted_Events_Run_1724
  • 32161 02/01/00 15:35 adev Accepted_Events_Run_1723
  • 32159 02/01/00 12:24 R. Michaels Trigger (and MLU) ok
  • 32156 02/01/00 12:14 adev Accepted_Events_Run_1721
  • 32153 02/01/00 11:10 R. Michaels possibly bad MLU
  • 32141 02/01/00 10:57 adev Accepted_Events_Run_1719
  • 32140 02/01/00 10:53 adev Accepted_Events_Run_1718
  • 32139 02/01/00 10:44 adev Accepted_Events_Run_1717
  • 32138 02/01/00 10:43 adev Accepted_Events_Run_1716
  • 32137 02/01/00 10:05 adev Accepted_Events_Run_1715
  • 32136 02/01/00 09:51 adev Accepted_Events_Run_1714
  • 32135 02/01/00 09:50 adev Accepted_Events_Run_1713
  • 32133 02/01/00 09:33 adev Accepted_Events_Run_1712
  • 32131 02/01/00 05:35 kuss Hall A in restricted access
  • 32130 02/01/00 04:28 kuss no beam tonight
  • 32129 02/01/00 03:37 kuss Magnet shed is open
  • 32128 02/01/00 02:21 kuss MCC status